### v1

# A Single-Loop First-Order Algorithm for Linearly Constrained Bilevel Optimization

Abstract
We study bilevel optimization problems where the lower-level problems are strongly convex and have coupled linear constraints. To overcome the potential non-smoothness of the hyper-objective and the computational challenges associated with the Hessian matrix, we utilize penalty and augmented Lagrangian methods to reformulate the original problem as a single-level one. Especially, we establish a strong theoretical connection between the reformulated function and the original hyper-objective by characterizing the closeness of their values and derivatives. Based on this reformulation, we propose a single-loop, first-order algorithm for linearly constrained bilevel optimization (SFLCB). We provide rigorous analyses of its non-asymptotic convergence rates, showing an improvement over prior double-loop algorithms – form $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$. The experiments corroborate our theoretical findings and demonstrate the practical efficiency of the proposed SFLCB algorithm. Simulation code is provided at https://github.com/ShenGroup/SFLCB.
1 Introduction
In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter, ; franceschi2018bilevel, ; mackay2019self, ), data hypercleaning (shaban2019truncated, ), meta learning (rajeswaran2019meta, ; ji2020convergence, ), reinforcement learning (sutton2018reinforcement, ; hong2023two, ) and neural architecture search (liu2018darts, ; liang2019darts+, ); see survey papers (zhang2024introduction, ; liu2021investigating, ; sinha2017review, ) for additional discussions. While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation, ; liu2020generic, ; ji2021bilevel, ; dagreou2022framework, ; hong2023two, ; kwon2023fully, ), studies focusing on constrained BLO problems are relatively limited.
In this paper, we consider the following BLO problem where the lower-level (LL) problem has coupled constraints:
$\min_{x\in\mathcal{X}}\quad\Phi(x)\triangleq f(x,y^{*}(x))$ (1)
$\text{s.t. }\quad y^{*}(x)\in\arg\min_{y\in\mathcal{Y}(x)}g(x,y).$
The upper-level (UL) objective function $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ and the lower-level objective function $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ are continuously differentiable. Moreover, we assume that $g(x,y)$ is strongly convex with respect to $y$. The feasible sets are defined as $\mathcal{X}=\mathbb{R}^{d_{x}},\mathcal{Y}(x)=\{\,y\in\mathbb{R}^{d_{y}}\mid h(x,y)\leq 0\,\}$ where $h:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{h}}$.
For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$. This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized, ), hyperparameter optimization of constrained learning problems (xu2023efficient, ) and adversarial training (zhang2022revisiting, ) and draw significant attentions (tsaknakis2022implicit, ; khanduri2023linearly, ; kornowski2024first, ).
A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ). For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ). However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ). The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
Recently, some first-order methods (kwon2023penalty, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ; kornowski2024first, ) have been proposed for addressing constrained BLO problems. Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem. For example, yao2024constrained ; yao2024overcoming reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem. However, neither yao2024constrained nor yao2024overcoming establishes clear relationships between the stationary points of their approximated problems and the original one.
Works most closely related to ours are those by kwon2023penalty and jiang2024primal , both of which reformulated the problem (1) as
$\min_{x\in\mathcal{X},y\in\mathcal{Y}(x)}f(x,y)\quad\text{s.t. }g(x,y)-\min_{z\in\mathcal{Y}(x)}g(x,z)\leq 0,$ (2)
and considered optimizing the following function with a penalty parameter $\delta$:
$\min_{x\in\mathcal{X}}[\Phi_{\delta}(x)\triangleq\min_{y\in\mathcal{Y}(x)}\max_{z\in\mathcal{Y}(x)}\Phi_{\delta}(x,y,z)]$ (3)
where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$. Based on this reformulation, kwon2023penalty proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and jiang2024primal proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. However, kwon2023penalty only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration. Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation. Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty, ; jiang2024primal, ) for coupled constraints $\mathcal{Y}(x)$.
To address these limitations, in this paper, we establish a rigorous theoretical justification for this reformulation (3) and propose a single-loop Hessian-free algorithm for the linearly constrained cases. Our main contributions can be summarized as follows.
  • We establish a rigorous theoretical connection between the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ by proving the closeness of their values and derivatives under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$ with certain assumptions, which provides strong justifications for the reformulation (3).
  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem. and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies). The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal .
  • Our experiments on hyperparameter optimization in the support vector machine (SVM) and transportation network design problems validate the practical effectiveness and efficiency of the proposed SFLCB algorithm.
Table 1: Comparison of our paper with (kwon2023penalty, ; jiang2024primal, ; kornowski2024first, ). More detailed introductions and discussions of other related works can be found in Section˜2. Here, the “Complexity” for (kwon2023penalty, ; jiang2024primal, ) and SFLCB means the complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$; for (kornowski2024first, ), it means the complexity needed to achieve the $(\delta,\epsilon)$-Goldstein stationarity defined in kornowski2024first .
Methods LL Constraint Complexity Loops
kwon2023penalty $y\in\mathcal{Y}$, $\mathcal{Y}$ is a convex and compact set $O(\epsilon^{-3}\log(\epsilon^{-1}))$ single/double
jiang2024primal $h(x,y)\leq 0$, LICQ holds $O(\epsilon^{-5}\log(\epsilon^{-1}))$ triple
jiang2024primal $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(d\delta^{-1}\epsilon^{-3})$ double
kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(\delta^{-1}\epsilon^{-4})$ double
SFLCB (ours) $Bx+Ay-b\leq 0$, $A$ is full row rank $O(\epsilon^{-3})$ single
SFLCB (ours) $Ay\leq 0$, LICQ holds at the initial points $O(\epsilon^{-3})$ single
SFLCB (ours) $Ay\leq 0$ $O(\epsilon^{-4})$ single
2 Related works
BLO without constraints. One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter, ). It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation, ). Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ; yang2023accelerating, ). Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient, ; grazzi2020iteration, ; liu2021towards, ; bolte2022automatic, ). Recently, penalty-based methods have gained traction as a promising approach for solving BLO. Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value, ; mehra2021penalty, ; liu2022bome, ; kwon2023fully, ; shen2023penalty, ; gao2022value, ; lu2024slm, ).
BLO with constraints. There are two primary types of methods for solving constrained bilevel optimization problems. One is based on the implicit gradient method. Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly, ). tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm. khanduri2023linearly proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$. xu2023efficient used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$. However, they only provided an asymptotic convergence analysis of their algorithm. xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem. However, these implicit gradient-type algorithms (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
Another commonly used approach for solving constrained BLO problems is based on penalty reformulation. For example, yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function. However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly. yao2024overcoming reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm. Compared to yao2024constrained , yao2024overcoming did not need the projection operator to the coupled constraint set. However, both yao2024constrained and yao2024overcoming did not establish very clear relationships between the stationary points of their approximated problems and the original one. For example, yao2024overcoming only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one. Recently, tsaknakis2023implicit ; jiang2024barrier proposed algorithms based on barrier approximation approach for constrained BLO problems. However, their algorithms also require the computations for the Hessian matrix.
kwon2023penalty and jiang2024primal considered the same reformulation as ours. kwon2023penalty studied the case where the LL variables $y\in\mathcal{Y}$ are independent of $x$ and characterized the conditions under which the values and derivatives of $\Phi$ and $\Phi_{\delta}$ can be $O(\delta)$-close for $y\in\mathcal{Y}$ constraints. Compared with kwon2023penalty , we prove similar results under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. Moreover, the algorithms in kwon2023penalty require the implementation of the projection operator to $\mathcal{Y}$ at each iteration, which can be costly. jiang2024primal studied the coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. While jiang2024primal considered more general constraints than ours, however, it did not characterize the gap between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$. Our Theorem˜4.9 provides further justifications for their reformulation in coupled constraints. Moreover, compared with the double- and triple-loop algorithms in kwon2023penalty , we propose a single-loop algorithm SFLCB and prove an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$.
Recently, kornowski2024first also proposed first-order methods for linearly constrained BLO. Especially, they proved a nearly optimal convergence rate $\tilde{O}(\epsilon^{-2})$ for linear equality constraints and proposed algorithms that can attain $(\delta,\epsilon)$-Goldstein stationarity for linear inequality constraints. However, their convergence rates for linear inequality constraints either have additional dependence on dimension $d$ (such as $\tilde{O}(d\delta^{-1}\epsilon^{-3})$) or need additional assumptions to access the exact optimal dual variable (such as $\tilde{O}(\delta^{-1}\epsilon^{-4})$), while we do not require the exact optimal dual variable assumption. Compared with the double-loop algorithms in kornowski2024first , our proposed single-loop one is easier to implement in practice. Moreover, our techniques are also different from theirs under linear inequality constraints, thereby highlighting the distinct contributions and independent interests of our work.
3 Preliminaries
Notation. For vectors $a,b\in\mathbb{R}^{d}$, we denote $a\leq b$ if for all $i\in[d]$, $a_{i}\leq b_{i}$. We use $\|\cdot\|$ to denote the $l_{2}$ norm of a vector and the spectral norm of a matrix. We define the projection operator that project $x$ to a set $\mathcal{P}$ as $\Pi_{\mathcal{P}}(x)=\mathop{\rm argmin}_{x^{\prime}\in\mathcal{P}}\frac{1}{2}\|x-x^{\prime}\|^{2}$. We denote the projection operator that projects a $x\in\mathbb{R}^{d}$ to the set $\mathbb{R}^{d}_{-}$ as $\Pi_{-}(x)$.
We state the following assumptions for problem (1), which are commonly used in the theoretical studies of BLO.
Assumption 3.1.
For any $x\in\mathcal{X}$, $\mathcal{Y}(x)$ is nonempty, closed, and convex and $\Phi(x)$ is lower bounded by a finite, $\Phi^{*}=\inf_{x\in\mathcal{X}}\Phi(x)\geq$-$\infty$.
Assumption 3.2.
$f,\nabla f,\nabla g$ are Lipschitz continuous with $l_{f,0},l_{f,1},l_{g,1}$ respectively, jointly over $\mathcal{X}\times\mathcal{Y}(x)$.
Assumption 3.3.
For any fixed $x\in\mathcal{X}$, $g(x,y)$ is $\mu_{g}$-strongly convex with respect to $y\in\mathcal{Y}(x)$.
We introduce the standard definition of the $\epsilon$-stationary point for a differentiable function.
Definition 3.4.
We say $\hat{x}$ is an $\epsilon$-stationary point of a differentiable function $f$ if $\|\nabla f(\hat{x})\|\leq\epsilon$.
4 Reformulation
In this section, we provide a theoretical justification for our reformulation (3) and establish the conditions under which the function values and gradients of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ become sufficiently close. Note that in this section, we considered general coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$ which include, but are not limited to, the linear constraint case. The complete proofs for the lemmas and theorems in this section can be found in Appendix˜C.
First, we assume $\delta\leq\frac{\mu_{g}}{2l_{f,1}}$ and introduce the following notations:
$y^{*}_{\delta}(x)=\mathop{\rm argmin}_{y\in\mathcal{Y}(x)}\delta f(x,y)+g(x,y)$
$y^{*}(x)=z^{*}(x)=\mathop{\rm argmin}_{y\in\mathcal{Y}(x)}g(x,y)$
$\phi_{\delta}(x,y,z)=\delta\Phi_{\delta}(x,y,z)$
$\phi_{\delta}(x)=\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x)).$
Similar to Theorem 3.8 in kwon2023penalty , we have the following theorem to bound the difference between $\Phi$ and $\Phi_{\delta}$, as well as $y^{*}(x)$ and $y_{\delta}^{*}(x)$ in the coupled constraints.
Theorem 4.1.
When ˜3.1, 3.2 and 3.3 hold, we have
$0\leq\Phi(x)-\Phi_{\delta}(x)\leq\frac{\delta l_{f,0}^{2}}{2\mu_{g}},\quad\|y_{\delta}^{*}(x)-y^{*}(x)\|\leq\frac{2\delta l_{f,0}}{\mu_{g}}.$
Theorem˜4.1 characterizes how the difference in function values and the optimal LL variables between the reformulated problem and the original one are controlled by the penalty parameter $\delta$. Therefore, by choosing a sufficiently small $\delta$, i.e., $\delta=O(\epsilon)$, we can treat the reformulated problem $\min_{x\in\mathcal{X}}\Phi_{\delta}(x)$ as an approximation of the original problem and solve this approximated problem instead. In the following lemmas, we will provide the conditions under which the reformulated function $\Phi_{\delta}(x)$ is differentiable. Before that, we first introduce the well-known and commonly used Linear Independence Constraint Qualification (LICQ) condition.
Definition 4.2 (Active set).
We denote $\mathcal{I}_{y}\subseteq[d_{h}]$ as the active set of $y$, i.e. $\mathcal{I}_{y}=\{i\in[d_{h}]\mid h_{i}(x,y)=0\}$.
Definition 4.3 (LICQ).
We say a point $y$ satisfy the LICQ condition if, for all $i\in\mathcal{I}_{y}$, $\nabla_{y}h_{i}(x,y)$ are linearly independent.
Then, similar to Lemmas 2 and 3 in jiang2024primal , we have the following lemma.
Lemma 4.4.
When ˜3.1, 3.2, 3.3 hold and $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, the LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, then there exist the corresponding unique Lagrangian multipliers $\lambda^{*}(x)\in\mathbb{R}^{d_{h}}$ and $\lambda^{*}_{\delta}(x)\in\mathbb{R}^{d_{h}}$ such that
$\lambda^{*}(x)=\mathop{\rm argmax}_{\lambda\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}g(x,y)+\lambda^{\top}h(x,y)$ (4)
$\lambda^{*}_{\delta}(x)=\mathop{\rm argmax}_{\lambda\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}\delta f(x,y)+g(x,y)+\lambda^{\top}h(x,y).$ (5)
Furthermore, we have
$\nabla\Phi_{\delta}(x)$ $=\nabla_{x}f(x,y^{*}_{\delta}(x))+\frac{1}{\delta}[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)$
$\qquad-\nabla_{x}g(x,y^{*}(x))-\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)].$
While the gradients of $\Phi_{\delta}(x)$ exists under LICQ conditions, for general problem (1), $\Phi(x)$ is not guaranteed to be differentiable. For example, khanduri2023linearly provides an example where the LICQ condition holds, $\Phi(x)$ is non-differentiable at some points. However, if a given $x$ satisfies the following conditions, then $\nabla\Phi(x)$ exists at $x$.
Assumption 4.5 (Strict Complementarity).
Let $\lambda^{*}(x)$ be the Lagrange multipliers for $y^{*}(x)$ (4). For any $i\in\mathcal{I}_{y^{*}(x)}$, $\lambda^{*}(x)]_{i}>0$.
Assumption 4.6.
$\nabla^{2}f,\nabla^{2}g$ are Lipschitz continuous with $l_{f,2},l_{g,2}$ respectively, jointly over $\mathcal{X}\times\mathcal{Y}(x)$. For $i\in[d_{h}]$, $h_{i}(x,y)$ is convex with respect to $y$, $h_{i},\nabla h_{i},\nabla^{2}h_{i}$ are respectively Lipschitz continuous with $l_{h,0},l_{h,1},l_{h,2}$ jointly over $\mathcal{X}\times\mathcal{Y}(x)$.
Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; jiang2024barrier, ; kwon2023penalty, ) to ensure the existence of $\nabla\Phi(x)$.
Lemma 4.7 (Theorem 2 in xu2023efficient ).
When ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and LICQ (Definition˜4.3) condition hold for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
Moreover, with additional assumptions, we can establish a non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
Assumption 4.8.
For any $t\in[0,\delta]$,
  1. (1)
    $y^{*}_{t}(x)$ satisfies the LICQ condition (Definition˜4.3) with the same active set as $y^{*}(x)$. Denote this active set as $\mathcal{I}$. Let $\lambda^{*}_{t}(x)$ be the Lagrange multiplier for $y^{*}_{t}(x)$ in (5). For any $i\in\mathcal{I}$, $\lambda^{*}_{t}(x)]_{i}>0$ (Strict Complementarity). We assume $\|\lambda^{*}_{t}(x)\|\leq\Lambda$, where $\Lambda$ is an $O(1)$ constant.
  2. (2)
    Denote $\nabla_{y}\bar{h}(x,y^{*}_{t}(x))=\nabla_{y}[h(x,y^{*}_{t}(x))]_{\mathcal{I}}$. The singular values of $\nabla_{y}\bar{h}(x,y^{*}_{t}(x))$ satisfy $\sigma_{\max}([\nabla_{y}\bar{h}(x,y^{*}_{t}(x)))\leq s_{\max}$, $\sigma_{\min}(\nabla_{y}\bar{h}(x,y^{*}_{t}(x)))\geq s_{\min}>0$, where $s_{\max},s_{\min}$ are $O(1)$ constants.
˜4.8 is made for $t\in[0,\delta]$. When $\delta$ is sufficiently small, i.e., $\delta=O(\epsilon)$, $y^{*}(x)$ and $y^{*}_{t}(x)$ are very close according to Theorem˜4.1. Thus, we expect that for $t\in[0,\delta]$, $y^{*}_{t}(x)$ will have similar properties as $y^{*}(x)$. Similar assumptions have also been used in kwon2023penalty to establish the non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
Theorem 4.9.
When ˜3.1, 3.2, 3.3, 4.6 hold and $\delta\leq\mu_{g}/(2l_{f,1})$, if ˜4.8 holds for a given $x$, we have
$\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|\leq O(\delta).$
Similar non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$ has been established in kwon2023penalty ; however, their bound is established only for the LL constraints $\mathcal{Y}$ that do not depend on $x$. Our Theorem˜4.9 provides a more general theoretical justification for the validity of the reformulation (3) for coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
5 The SFLCB Algorithm
In the last section, we have justified the validity of our reformulation for coupled constrained BLO. In this section, we focus on a special and important case where the LL constraints are h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b. This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized, ; khanduri2023linearly, ), adversarial training (zhang2022revisiting, ; khanduri2023linearly, ), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6). For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty, ; jiang2024primal, ).
First, we introduce the following slackness parameters $\alpha,\beta\in\mathbb{R}^{d_{h}}_{-}$ and define $y^{\prime}=(y^{\top},\alpha^{\top})^{\top}$, $z^{\prime}=(z^{\top},\beta^{\top})^{\top}$. With these slackness parameters, we can convert the original inequality constraints to equality constraints, i.e. we can reformulate $\min_{x\in\mathcal{X},y\in\mathcal{Y}(x)}\max_{z\in\mathcal{Y}(x)}\phi_{\delta}(x,y,z)$ as:
$\min_{x\in\mathcal{X},y^{\prime}\in\mathcal{S}_{y}(x)}\max_{z^{\prime}\in\mathcal{S}_{y}(x)}\phi_{\delta}(x,y,z)$ (6)
where $\mathcal{P}_{y}=\{y\in\mathbb{R}^{d_{y}},\alpha\in\mathbb{R}^{d_{h}}_{-}\}$, $\mathcal{S}_{y}(x)=\{y,\alpha\in\mathcal{P}_{y}|h(x,y)-\alpha=0\}$. The Lagrangian of (6) with multiplier $u,v\in\mathbb{R}^{d_{h}}$ is
$L_{\delta}(x,y^{\prime},z^{\prime},u,v)=$ $\phi_{\delta}(x,y,z)+u^{\top}(h(x,y)-\alpha)-v^{\top}(h(x,z)-\beta).$
According to Proposition 5.3.4 in bertsekas2009convex , we know that
$\phi_{\delta}(x)=\min_{y^{\prime}\in\mathcal{P}_{y},v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y},u\in\mathbb{R}^{d_{h}}}L_{\delta}(x,y^{\prime},z^{\prime},u,v).$
Note that when $\delta\leq\mu_{g}/(2l_{f,1})$, $\phi_{\delta}$ is $\mu_{g}/2$-strongly convex with respect $y$. However, $L_{\delta}(x,y^{\prime},z^{\prime},u,v)$ is only convex with respect $y^{\prime}$ and concave with respect $z^{\prime}$. To make the objective function strongly convex with respect to $y^{\prime}$ and strongly concave with respect to $z^{\prime}$, we can construct an augmented Lagrangian $K$:
$K(x,y^{\prime},z^{\prime},u,v)=$ $L_{\delta}(x,y^{\prime},z^{\prime},u,v)+\frac{\rho_{1}}{2}\|h(x,y)-\alpha\|^{2}-\frac{\rho_{2}}{2}\|h(x,z)-\beta\|^{2}.$
With $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$ and $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, according to Lemma˜D.1, $K$ is strongly convex with respect to $y^{\prime}$ and strongly concave with respect to $z^{\prime}$. Moreover, we have
$\min_{y^{\prime}\in\mathcal{P}_{y},v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y},u\in\mathbb{R}^{d_{h}}}L_{\delta}(x,y^{\prime},z^{\prime},u,v)=\min_{y^{\prime}\in\mathcal{P}_{y},v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y},u\in\mathbb{R}^{d_{h}}}K(x,y^{\prime},z^{\prime},u,v).$
Note that $L_{\delta}$ and $K$ have the same optimal points and same optimal function value. Thus, we can reformulate the problem (6) to the minimax optimization problem over $K$:
$\min_{x\in\mathcal{X},y^{\prime}\in\mathcal{P}_{y},v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y},u\in\mathbb{R}^{d_{h}}}K(x,y^{\prime},z^{\prime},u,v).$ (7)
Motivated by these theoretical analyses, and applying gradient descent ascent (GDA) over problem (7), we propose SFLCB. A compact description can be found in Algorithm˜1.
Algorithm 1 SFLCB
Input: $\delta$, $\rho_{1}$, $\rho_{2}$, $\eta_{x},\eta_{y},\eta_{z},\eta_{v},\eta_{u},T$
Initialize: $x_{0}\in\mathcal{X},y^{\prime}_{0},z^{\prime}_{0}\in\mathcal{P}_{y},u_{0},v_{0}\in\mathbb{R}^{d_{h}}$
for t=0,1,,T1t=0,1,...,T-1 do
  $u_{t+1}=u_{t}+\eta_{u}(h(x_{t},y_{t})-\alpha_{t})$
  $v_{t+1}=v_{t}+\eta_{v}(h(x_{t},z_{t})-\beta_{t})$
  $x_{t+1}=x_{t}-\eta_{x}\nabla_{x}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})$
  $y^{\prime}_{t+1}=\Pi_{\mathcal{P}_{y}}\{y^{\prime}_{t}-\eta_{y}\nabla_{y}^{\prime}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})\}$
  $z^{\prime}_{t+1}=\Pi_{\mathcal{P}_{y}}\{z^{\prime}_{t}+\eta_{z}\nabla_{z}^{\prime}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})\}$
end for
5.1 Convergence results
In this section, we provide the non-asymptotic convergence results of SFLCB (Algorithm˜1) for two constraint settings: 1) h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $A$ is full row rank, and 2) h(y)=Aybh(y)=Ay-b, where $A$ is not required to be full row rank.
Note that when the LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, the optimal Lagrangian multipliers of $y^{*}(x)$ and $y^{*}_{\delta}(x)$ are unique. Thus, we first introduce the following lemma and notations for these optimal Lagrangian multipliers.
Lemma 5.1.
When the LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, the optimal Lagrangian multipliers of $y^{*}(x)$ and $y^{*}_{\delta}(x)$ are unique, and we have
$u^{*}_{\delta}(x)$ $=\mathop{\rm argmax}_{u\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)+u^{\top}h(x,y)=\mathop{\rm argmax}_{u\in\mathbb{R}^{d_{h}}}\min_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v),$
$v^{*}(x)$ $=\mathop{\rm argmax}_{v\in\mathbb{R}_{+}}\min_{z\in\mathcal{Y}(x)}g(x,z)+v^{\top}h(x,z)=\mathop{\rm argmin}_{v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v).$
The proof of Lemma˜5.1 can be found in Appendix˜E.
Next, we introduce the following notations:
$y^{\prime*}_{\delta}(x,u)=\mathop{\rm argmin}_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v),\quad z^{\prime*}(x,v)=\mathop{\rm argmax}_{z^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v),\quad A^{\prime}=(A,-I).$
Then, we present the convergence results of SFLCB (Algorithm˜1) for coupled constraints h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $A$ is full row rank. Note that BLOCC in jiang2024primal that achieves the complexity of $O(\epsilon^{-3}\log(\epsilon^{-1}))$ also needs the matrix $A$ to be full row rank (See Table˜1). For $A$ that is not full row rank and $B=0$, we provide the convergence results in Theorem˜5.4 and Corollary˜5.5.
Theorem 5.2.
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank, ˜3.1, 3.2, 3.3 hold and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply Algorithm˜1 with appropriate parameters (see Appendix˜E), then we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-4})$.
Moreover, if we have initial points $x_{0},y_{0},z_{0},u_{0},v_{0}$ such that
$\|y_{0}-y^{*}_{\delta}(x_{0})\|\leq O(\delta),\,\|u_{0}-u^{*}_{\delta}(x_{0})\|\leq O(\delta),$ (8)
$\|z_{0}-z^{*}(x_{0})\|\leq O(\delta),\,\|v_{0}-v^{*}(x_{0})\|\leq O(\delta),$ (9)
then we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$.
The formal statement and the complete proof of Theorem˜5.2 can be found in Appendix˜E.
Proof sketch for Theorem˜5.2. The key new idea in our proof is the construction of a novel potential function $V_{t}$ and prove the descent lemma of $V_{t}$ (Lemma˜E.4). $V_{t}$ is defined as:
$V_{t}=\frac{1}{4}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{t},v_{t})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
where $d(x,z^{\prime},u,v)=K(x,y^{\prime*}_{\delta}(x,u),z^{\prime},u,v)$ and $q(x,v)=\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x,v))-v^{\top}(A^{\prime}z^{\prime*}(x,v)-b)-\frac{\rho_{2}}{2}\|A^{\prime}z^{\prime*}(x,v)-b\|^{2}$. To prove Lemma˜E.4, we need to first prove several novel error bounds in Lemma˜D.2, Lemma˜E.3 and Lemma˜E.2. Those error bounds may be of independent interest for solving other similar problems. The full row rank property of $A$ is used in Lemma˜E.2 to bound $\|u_{t+1}-u^{*}_{\delta}(x_{t})\|$ and $\|v_{t+1}-v^{*}(x_{t})\|$.
Since $A$ has full row rank, then according to Theorem 6 in jiang2024primal , we can easily find initial points satisfying (8)-(9) with a complexity of $O(\log(\epsilon^{-1}))$ and we have the following corollary.
Corollary 5.3.
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA has full row rank, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply projected gradient descent (PGD) for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, we can find $\hat{v}$, $\hat{z}$ such that $\|\hat{v}-v^{*}(x_{0})\|\leq\delta$ and $\|\hat{z}-z^{*}(x_{0})\|\leq\delta$ with a complexity of $O(\log(\epsilon^{-1}))$. Set $y_{0}=z_{0}=\hat{z}$, $u_{0}=v_{0}=\hat{v}$, $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$. With $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points and applying Algorithm˜1, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$. Thus, the total complexity is $O(\epsilon^{-3}+\log(\epsilon^{-1}))=O(\epsilon^{-3})$.
The proof of Corollary˜5.3 can be found in Appendix˜E. Thus, compared to jiang2024primal , we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ for the coupled linear constraint (See Table˜1).
Additionally, we have the following convergence results for constraints h(y)=Aybh(y)=Ay-b, where $A$ is not required to have a full row rank.
Theorem 5.4.
When h(x,y)=Aybh(x,y)=Ay-b, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply Algorithm˜1 with appropriate parameters (see Appendix˜D), then we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-4})$.
Moreover, if we have initial points $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ such that
$\|y_{0}-y^{*}_{\delta}(x_{0})\|\leq O(\delta),\,\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|\leq O(\delta),\,\|A^{\prime}y_{0}^{\prime}-b\|\leq O(\delta)$ (10)
$\|z_{0}-z^{*}(x_{0})\|\leq O(\delta),\,\|A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|\leq O(\delta)\,\|A^{\prime}z_{0}^{\prime}-b\|\leq O(\delta)$ (11)
then we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$.
The formal statement and the complete proof of Theorem˜5.4 can be found in Appendix˜D.
Proof sketch for Theorem˜5.4. The general proof flow of Theorem˜5.4 is similar to that of Theorem˜5.2. However, since here we do not have coupled constraints, $\nabla_{x}K$ has no relationship to $u$ or $v$, and according to the Danskin’s theorem, $\nabla\phi_{\delta}(x)=\delta\nabla_{x}f(x,y^{*}_{\delta}(x))+\nabla_{x}g(x,y^{*}_{\delta}(x))-\nabla_{x}g(x,z^{*}(x))$ also has no relationship to $u^{*}_{\delta}(x)$ or $v^{*}(x)$. Thus, we do not require Lemma˜E.2 or the full row-rank assumption on $A$ in this setting.
Next, we show that, as long as the LICQ condition (Definition˜4.3) holds for the initial $y^{*}(x_{0})$ and $y^{*}_{\delta}(x_{0})$, we can find initial points satisfying (10)-(11) with a complexity of $O(\epsilon^{-2})$ and we have the following corollary.
Corollary 5.5.
When h(x,y)=Aybh(x,y)=Ay-b, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, for a given initial point $x_{0}$, if the LICQ condition (Definition˜4.3) holds at $y^{*}(x_{0})$ and $y^{*}_{\delta}(x_{0})$, we can apply Algorithm˜1 with fixed $x_{0}$. Then for a sufficiently small $\epsilon$ (see Appendix˜D), we can find $\hat{y}^{\prime},\hat{z}^{\prime},\hat{u},\hat{v}$ such that
$\|\hat{y}-y^{*}_{\delta}(x_{0})\|\leq O(\delta),\,\|A^{\prime}y^{\prime*}_{\delta}(x_{0},\hat{u})-b\|\leq O(\delta),\,\|A^{\prime}\hat{y}^{\prime}-b\|\leq O(\delta),\,\|\hat{u}-u^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|\hat{z}-z^{*}(x_{0})\|\leq O(\delta),\,\|A^{\prime}z^{\prime*}(x_{0},\hat{v})-b\|\leq O(\delta),\,\|A^{\prime}\hat{z}^{\prime}-b\|\leq O(\delta),\,\|\hat{v}-v^{*}(x_{0})\|\leq O(\delta)$
with a complexity of $O(\epsilon^{-2})$. Set $y^{\prime}_{0}=\hat{y}^{\prime}$, $z^{\prime}_{0}=\hat{z}^{\prime}$, $u_{0}=\hat{u}$, $v_{0}=\hat{v}$. With $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$. Thus, the total complexity is $O(\epsilon^{-3}+\epsilon^{-2})=O(\epsilon^{-3})$.
The proof of Corollary˜5.5 can be found in Appendix˜D. The key to proving Corollary˜5.5 lies in Lemma˜D.4. In Lemma˜D.4, we prove that, without the full row-rank assumption on $A$, we can bound $\|u_{t+1}-u^{*}_{\delta}(x_{0})\|$ and $\|v_{t+1}-v^{*}(x_{0})\|$ with a fixed $x_{0}$. Thus, we can use our algorithm SFLCB to find suitable initial points with a fixed $x_{0}$.
Note that in Corollary˜5.5, we only need the LICQ condition holds for the initial $y^{*}(x_{0})$ and $y^{*}_{\delta}(x_{0})$, and we can achieve a total complexity of $O(\epsilon^{-3})$. Compared to the decoupled constrained setting in kwon2023penalty , we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ (See Table˜1).
6 Experiments
In this section, we evaluate the performance of our SFLCB algorithm on three tasks: a toy example, hyperparameter optimization for SVM, and a transportation network design problem. These experiments demonstrate the practical effectiveness and efficiency of SFLCB. Additional hype-parameter sensitivity analysis experiments and detailed experimental settings can be found in Appendix˜F.
6.1 Toy example
Figure 1: Toy example.
We consider the same constrained BLO problem that was studied in jiang2024primal , which is:
$\min_{x\in[0,3]}f(x,y^{*}(x))=\frac{e^{-y_{g}^{*}(x)+2}}{2+\cos(6x)}+\frac{1}{2}\ln\left((4x-2)^{2}+1\right)$ (12)
$\text{s.t. }\quad y^{*}(x)\in\arg\min_{y\in\mathcal{Y}(x)}g(x,y)=(y-2x)^{2}$ (13)
where $\mathcal{Y}(x)=\{y\in\mathbb{R}|y\leq x\}$. Note that this problem satisfies ˜3.1, 3.2, 3.3. Moreover, we have $y^{*}(x)=x$ and Equation˜12 is equivalent to $\min_{x\in[0,3]}f(x,y)|_{y=x}$. In Figure˜1, we plot the hyper-objective function $f(x,y^{*}(x))$. The red points indicate the converged solutions obtained by our algorithm with 200 different initialization values. We notice that SFLCB consistently finds the local minima of the hyper-objective function, which validates the effectiveness of SFLCB.
Figure 2: Hyperparameter optimization in SVM.
6.2 Hyperparameter optimization in SVM
Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ). Here we consider the same problem formulation as in jiang2024primal , which formulates this problem as a coupled linearly constrained BLO problem. We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient, ), LV-HBA (yao2024constrained, ), BLOCC (jiang2024primal, ), and BiC-GAFFA (yao2024overcoming, ) on the diabetes dataset (dua2017uci, ). Results are plotted in Figure˜2. We notice that our SFLCB algorithm converges significantly faster than other algorithms, which demonstrates the practical efficiency of the proposed SFLCB algorithm.
6.3 Transportation network design
We further conduct experiments on a transportation network design problem, following the same setting as in jiang2024primal . In this setting, we act as the operator, whose profit serves as the upper-level objective and is influenced by passenger behavior, which is modeled in the lower-level problem. Detailed formulations and settings can be found in Appendix˜F. We consider the two synthetic networks of 3 and 9 nodes, same as those considered in jiang2024primal . We compare SFLCB with BLOCC (jiang2024primal, ). Results are plotted in Figure˜3, which indicate that SFLCB significantly outperforms BLOCC on this network design task.
(a) 3 nodes
(b) 9 nodes
Figure 3: Results of the transportation experiments on 3 nodes and 9 nodes settings. Larger UL utility indicates better performance.
6.3.1 Sensitivity analysis of $\delta$
We also conduct the sensitivity analysis of the $\delta$ in SFLCB for the 3-node network. We set $\rho_{1}=\rho_{2}=1000$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-4$, and $T=20000$. Then, we test different $\delta$ values from $\{0.01,0.05,0.1,0.5,1\}$. For each $\delta$, we test with three different random seeds. The final average results and one standard deviation are reported in Figure˜4. As can be seen, larger values of $\delta$ lead to a faster initial decrease in the loss (increase in UL utility). In contrast, very small $\delta$ (e.g., $\delta=0.01$) results in significantly slower convergence overall. However, an overly large $\delta$ (e.g., $\delta=1$) can lead to large approximation errors in later stages, causing deviation from the true optimization objective and ultimately poor performance. We observe that moderate values of $\delta$ (such as 0.05, 0.1, and 0.5) achieve relatively good final performance. These observations are consistent with our theoretical predictions. For example, our theory indicates that the convergence rate of SLFCB is inversely proportional to $\delta$: smaller $\delta$ leads to slower convergence but smaller approximation error, whereas larger $\delta$ improves convergence speed towards the approximate problem but incurs greater approximation error. Figure˜4 indicates a properly chosen $\delta$ thus can balance convergence speed and approximation error.
Figure 4: Comparison of different $\delta$ in SFLCB for the 3 nodes network.
7 Conclusions and future directions
In this paper, for coupled constrained BLO problem in Equation˜1, we theoretically analyzed the relationship between the original hyper-objective $\Phi$ and the reformulated function $\Phi_{\delta}$ in Equation˜3, providing a solid justification for the validity of the reformulation. Especially, for the linearly constrained case, we proposed SFLCB, a single-loop, Hessian-free algorithm, improving the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ over previous works kwon2023penalty ; jiang2024primal . Our experiments on hyperparameter optimization for SVM and the transportation network design problem validated the practical efficiency of the proposed SFLCB algorithm. One limitation of our work is that the analysis is restricted to deterministic and linearly constrained settings. A promising direction for future research is to extend the current results to stochastic environments or more general constraint structures. Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal, ) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.
Acknowledgements
The work of Wei Shen and Cong Shen was supported in part by the US National Science Foundation (NSF) under awards 2143559 and 2332060. The work of Jiawei Zhang was supported by the startup funding from the University of Wisconsin-Madison.
Appendix
The Appendix is organized as follows. In Appendix˜A, we introduce some useful lemmas that will be utilized in the subsequent proofs. In Appendix˜B, we introduce some notations that will be used throughout the Appendix. In Appendix˜C, we provide the proofs for the lemmas and theorems in Section˜4. In Appendix˜D, we provide proofs for Theorem˜5.4 and Corollary˜5.5. In Appendix˜E, we provide proofs for Theorem˜5.2 and Corollary˜5.3. In Appendix˜F, we present the detailed experimental settings along with additional hyperparameter sensitivity analysis experiments.
Appendix A Useful Lemmas
Lemma A.1 (Lemma 12 in [3]).
Suppose $f(\cdot)$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{X}$ is a convex closed set, and $\eta\leq 1/l$. Define $x^{*}=\mathop{\rm argmin}_{x\in\mathcal{X}}f(x)$ and $x^{+}=\Pi_{\mathcal{X}}(x-\eta\nabla f(x))$. Then, we have
$\|x-x^{*}\|\leq\frac{2}{\mu\eta}\|x-x^{+}\|$
Lemma A.2 (Lemma 23 in [24]).
Suppose for any fixed $y\in\mathcal{Y}$, $f(x,y)$ is $\mu$-strongly convex w.r.t. $x$ and suppose for any $y_{1},y_{2}\in\mathcal{Y},x\in\mathcal{X}$, $\|\nabla_{x}f(x,y_{1})-\nabla_{x}f(x,y_{2})\|\leq l\|y_{1}-y_{2}\|$. Define $x^{*}(y)=\mathop{\rm argmin}_{x\in\mathcal{X}}f(x,y)$. Then, we have
$\|x^{*}(y_{1})-x^{*}(y_{2})\|\leq\frac{l}{\mu}\|y_{1}-y_{2}\|$
Lemma A.3 (Theorem 4.1 in [50]).
Suppose $f(x):\mathbb{R}^{d_{1}}\to\mathbb{R}$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{P}=\{x|Cx\leq e\},\mathcal{S}(r)=\{x\in\mathcal{P}|Ax-b=r\}$, where $C,A\in\mathbb{R}^{d_{2}\times d_{1}}$, $e,b,r\in\mathbb{R}^{d_{2}}$. Define
$L(x,u)=f(x)+u^{\top}(Ax-b)$
$x^{*}(u)=\mathop{\rm argmin}_{x\in\mathcal{P}}L(x,u)$
$x^{*}_{r}=\mathop{\rm argmin}_{x\in\mathcal{S}(r)}f(x)$
where $u\in\mathbb{R}^{d_{2}}$ is the Lagrange multiplier. We have
$\|x^{*}(u)-x^{*}_{0}\|\leq\sigma_{x}\|Ax^{*}(u)-b\|$
$\|x^{*}_{r}-x^{*}_{0}\|\leq\sigma_{x}\|r\|$
where
$\sigma_{x}=\frac{\sqrt{2}(\bar{\theta}l^{2}+1)}{\mu}$
$M=\begin{pmatrix}A^{\top}&C^{\top}\\ 0&I\end{pmatrix}$
$\bar{\theta}=\max_{\bar{M}\in\mathcal{B}(M)}\sigma^{2}_{\max}(\bar{M})/\sigma^{4}_{\min}(\bar{M})$
$\bar{M}$ is the set of all submatrices of $M$ with full row rank.
Lemma A.4.
Suppose $f(x):\mathbb{R}^{d_{x}}\to\mathbb{R}$ is $\mu$-strongly convex w.r.t. $x$. Define $L(x,\alpha,u)=f(x)+u^{\top}(Ax-b-\alpha)+\frac{\rho}{2}\|Ax-b-\alpha\|^{2}$, where $\alpha,u,b\in\mathbb{R}^{d_{y}}$, $A\in\mathbb{R}^{d_{y}\times d_{x}}$, $\rho\in(0,\mu/\sigma^{2}_{\max}(A))$. Denoting $x^{\prime}=(x^{\top},\alpha^{\top})^{\top}$, $L(x^{\prime},u)=L(x,\alpha,u)$, we have $L(x^{\prime},u)$ is $\mu_{x}$-strongly convex w.r.t. $x^{\prime}$, where $\mu_{x}=\min\{\mu-\rho\sigma_{\max}^{2}(A),\frac{\rho}{2}\}$.
Proof.
For any $x_{1},x_{2}\in\mathbb{R}^{d_{x}},\alpha_{1},\alpha_{2}\in\mathbb{R}^{d_{y}}$, denoting $x^{\prime}_{1}=(x_{1}^{\top},\alpha_{1}^{\top})^{\top}$, $x^{\prime}_{2}=(x_{2}^{\top},\alpha_{2}^{\top})^{\top}$, $A^{\prime}=(A,-I)$, we have
$L(x^{\prime}_{1},u)-L(x^{\prime}_{2},u)-\langle\nabla_{x^{\prime}}L(x^{\prime}_{2},u),x^{\prime}_{1}-x^{\prime}_{2}\rangle$
$=$ $f(x_{1})-f(x_{2})+u^{\top}A^{\prime}(x^{\prime}_{1}-x^{\prime}_{2})+\frac{\rho}{2}\|A^{\prime}x^{\prime}_{1}-b\|^{2}-\frac{\rho}{2}\|A^{\prime}x^{\prime}_{2}-b\|^{2}$
$-\langle\nabla_{x}f(x_{2}),x_{1}-x_{2}\rangle-u^{\top}A^{\prime}(x^{\prime}_{1}-x^{\prime}_{2})-\rho(A^{\prime\top}A^{\prime}x^{\prime}_{2}-A^{\prime\top}b)^{\top}(x^{\prime}_{1}-x^{\prime}_{2})$
$\geq$ $\frac{\mu}{2}\|x_{1}-x_{2}\|^{2}+\frac{\rho}{2}\|A^{\prime}(x^{\prime}_{1}-x^{\prime}_{2})\|^{2}$
$\geq$ $\frac{\mu}{2}\|x_{1}-x_{2}\|^{2}+\frac{\rho}{4}\|\alpha_{1}-\alpha_{2}\|^{2}-\frac{\rho}{2}\|A(x_{1}-x_{2})\|^{2}$
$\geq$ $\frac{\mu-\rho\sigma_{\max}^{2}(A)}{2}\|x_{1}-x_{2}\|^{2}+\frac{\rho}{4}\|\alpha_{1}-\alpha_{2}\|^{2}$
$\geq$ $\frac{\mu_{x}}{2}\|x^{\prime}_{1}-x^{\prime}_{2}\|^{2}$
where $\mu_{x}=\min\{\mu-\rho\sigma_{\max}^{2}(A),\frac{\rho}{2}\}$. ∎
Appendix B Notations
Denote $l_{\delta}=\mu_{g}/2$. We introduce the following notations.
$\Phi_{\delta}(x,y,z)=\frac{1}{\delta}\left(\delta f(x,y)+g(x,y)-g(x,z)\right)$
$\phi_{\delta}(x,y,z)=\delta f(x,y)+g(x,y)-g(x,z)$
$g_{\delta}(x,y)=\delta f(x,y)+g(x,y)$
$y^{*}_{\delta}(x)=\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)$
$y^{*}(x)=\min_{y\in\mathcal{Y}(x)}g(x,y)$
$L_{g}=l_{\delta}+l_{g,1}$
$L_{\phi}=l_{\delta}+2l_{g,1}$
$\Phi(x)=f(x,y^{*}(x))$
$\Phi_{\delta}(x)=\Phi_{\delta}(x,y^{*}_{\delta}(x),y^{*}(x))$
$\phi_{\delta}(x)=\phi_{\delta}(x,y^{*}_{\delta}(x),y^{*}(x))$
When $\delta\leq\mu_{g}/2l_{f,1}$, we have $\delta l_{f,1}\leq\mu_{g}/2=l_{\delta}$ and thus, $g_{\delta}(x,y)$ is $L_{g}$-smooth and $\phi_{\delta}(x,y,z)$ is $L_{\phi}$-smooth.
Appendix C Reformulation
In this section, we provide the proofs for the lemmas and theorems in Section˜4.
Theorem˜4.1
When ˜3.1, 3.2 and 3.3 hold, we have
$0\leq\Phi(x)-\Phi_{\delta}(x)\leq\frac{\delta l_{f,0}^{2}}{2\mu_{g}}$
$\|y_{\delta}^{*}(x)-y^{*}(x)\|\leq\frac{2\delta l_{f,0}}{\mu_{g}}.$
Proof.
Similar results and proofs of Theorem˜4.1 can also be found in [21, 16]. For completeness, we also provide our proofs for Theorem˜4.1 here.
Note that $y_{\delta}^{*}(x)$ satisfy
$\Pi_{\mathcal{Y}(x)}(y_{\delta}^{*}(x)-[\delta\nabla_{y}f(x,y_{\delta}^{*}(x))+\nabla_{y}g(x,y_{\delta}^{*}(x))]/L_{\phi})=y_{\delta}^{*}(x)$
Since $g(x,\cdot)$ is $\mu_{g}$-strongly concave and $l_{g,1}$-smooth and $L_{\phi}\geq l_{g,1}$, according to Lemma˜A.1, we have
$\|y_{\delta}^{*}(x)-z^{*}(x)\|\leq$ $\frac{2L_{\phi}}{\mu_{g}}\|y_{\delta}^{*}(x)-\Pi_{\mathcal{Y}(x)}(y_{\delta}^{*}(x)-\nabla_{y}g(x,y_{\delta}^{*}(x))/L_{\phi})\|$
$=$ $\frac{2L_{\phi}}{\mu_{g}}\|\Pi_{\mathcal{Y}(x)}(y_{\delta}^{*}(x)-[\delta\nabla_{y}f(x,y_{\delta}^{*}(x))+\nabla_{y}g(x,y_{\delta}^{*}(x))]/L_{\phi})$
$-\Pi_{\mathcal{Y}(x)}(y_{\delta}^{*}(x)-\nabla_{y}g(x,y_{\delta}^{*}(x))/L_{\phi})\|$
$\leq$ $\frac{2\delta}{\mu_{g}}\|\nabla_{y}f(x,y_{\delta}^{*}(x))\|$
$\leq$ $\frac{2\delta l_{f,0}}{\mu_{g}}.$
For $\Phi_{\delta}(x)$, we have
$\Phi_{\delta}(x)=$ $f(x,y_{\delta}^{*}(x))+\frac{1}{\delta}(g(x,y_{\delta}^{*}(x))-g(x,y^{*}(x)))$
$\leq$ $f(x,y^{*}(x))+\frac{1}{\delta}(g(x,y^{*}(x))-g(x,y^{*}(x)))$
$=$ $\Phi(x),$
and
$\Phi_{\delta}(x)=$ $f(x,y_{\delta}^{*}(x))+\frac{1}{\delta}[g(x,y^{*}_{\delta}(x))-g(x,y^{*}(x))]$
$\geq$ $f(x,y_{\delta}^{*}(x))+\frac{\mu_{g}}{2\delta}\|y_{\delta}^{*}(x)-y^{*}(x)\|^{2}$
$\geq$ $f(x,y^{*}(x))+\frac{\mu_{g}}{2\delta}\|y_{\delta}^{*}(x)-y^{*}(x)\|^{2}-l_{f,0}\|y_{\delta}^{*}(x)-y^{*}(x)\|$
$\geq$ $\Phi(x)-\frac{\delta l_{f,0}^{2}}{2\mu_{g}},$
where the first equality is due to the quadratic growth of a strongly convex function, the last equality is due to $ax^{2}+bx\geq-b^{2}/(4a)$.
Lemma˜4.4
When ˜3.1, 3.2, 3.3 hold and $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) hold for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, then there exist the corresponding unique Lagrangian multipliers $\lambda^{*}(x)\in\mathbb{R}^{d_{h}}$ and $\lambda^{*}_{\delta}(x)\in\mathbb{R}^{d_{h}}$ such that
$\lambda^{*}(x)=\mathop{\rm argmax}_{\lambda\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}g(x,y)+\lambda^{\top}h(x,y)$ (14)
$\lambda^{*}_{\delta}(x)=\mathop{\rm argmax}_{\lambda\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}\delta f(x,y)+g(x,y)+\lambda^{\top}h(x,y).$ (15)
Furthermore, we have
$\nabla\Phi_{\delta}(x)=$ $\nabla_{x}f(x,y^{*}_{\delta}(x))+\frac{1}{\delta}[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)$
$-\nabla_{x}g(x,y^{*}(x))-\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)].$ (16)
Proof.
The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [43]. According to Lemmas 2 and 3 in [16]. We know that when ˜3.1, 3.2, 3.3 hold, $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, defining $\psi(x)=g(x,y^{*}(x))$ and $\psi_{\delta}(x)=g_{\delta}(x,y^{*}_{\delta}(x))$, we have
$\nabla_{x}\psi(x)=\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)$
$\nabla_{x}\psi_{\delta}(x)=\delta\nabla_{x}f(x,y^{*}_{\delta}(x))+\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)$
Since $\Phi_{\delta}(x)=f(x,y_{\delta}^{*}(x))+\frac{1}{\delta}(g_{\delta}(x,y^{*}_{\delta}(x))-g(x,y^{*}(x)))$, we have
$\nabla\Phi_{\delta}(x)=$ $\nabla_{x}f(x,y^{*}_{\delta}(x))+\frac{1}{\delta}[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)$
$-\nabla_{x}g(x,y^{*}(x))-\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)].$
Lemma˜4.7
When ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ condition (Definition˜4.3) holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$ and can be expressed as
$\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$
where $\nabla y^{*}(x)$ can be calculated according to (20).
Proof.
According to Theorem 2 in [45], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
Moreover, we can give the explicit expression of $\nabla\Phi(x)$.
With $\lambda\in\mathbb{R}^{d_{h}}$, we have the following Lagrangian function
$\mathcal{L}(x,y,\lambda)=g(x,y)+\lambda^{\top}h(x,y)$
Denote $\lambda^{*}(x)$ as the optimal Lagrangian multiplier, $\mathcal{I}_{x}\subseteq[d_{h}]$ as the active set of $y^{*}(x)$, i.e. $\mathcal{I}_{x}=\{i\in[d_{h}][h(x,y)]_{i}=0\}$ Denote $\bar{h}(x,y^{*}(x))=[h(x,y^{*}(x))]_{\mathcal{I}_{x}}$ and $\bar{\lambda}^{*}(x)=[\lambda^{*}(x)]_{\mathcal{I}_{x}}$. We have the following KKT conditions:
$\nabla_{y}g(x,y^{*}(x))+\nabla_{y}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)=0$
$\bar{h}(x,y^{*}(x))=0.$
Differentiating the KKT conditions with respect to $x$, we have
$\nabla^{2}_{xy}g(x,y^{*}(x))+\nabla^{2}_{yy}g(x,y^{*}(x))\nabla y^{*}(x)+\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)$
$+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\nabla y^{*}(x)+\nabla_{y}\bar{h}(x,y^{*}(x))^{\top}\nabla\bar{\lambda}^{*}(x)=0$
$\nabla_{x}\bar{h}(x,y^{*}(x))+\nabla_{y}\bar{h}(x,y^{*}(x))\nabla y^{*}(x)=0$
Thus, $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ satisfy the following equation.
$\begin{bmatrix}\nabla^{2}_{yy}g(x,y^{*}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)&\nabla_{y}\bar{h}(x,y^{*}(x))^{\top}\\ \nabla_{y}\bar{h}(x,y^{*}(x))&0\end{bmatrix}\begin{bmatrix}\nabla y^{*}(x)\\ \nabla\bar{\lambda}^{*}(x)\end{bmatrix}$ (17)
$=$ $\begin{bmatrix}-\nabla^{2}_{xy}g(x,y^{*}(x))-\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ -\nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}$ (18)
Denote
$H=\begin{bmatrix}\nabla^{2}_{yy}g(x,y^{*}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)&\nabla_{y}\bar{h}(x,y^{*}(x))^{\top}\\ \nabla_{y}\bar{h}(x,y^{*}(x))&0\end{bmatrix}$ (19)
Since $g$ is strongly convex, $h$ is convex, $\nabla^{2}_{yy}g(x,y^{*}(x))\succ 0$, $\nabla^{2}_{yy}[\bar{h}(x,y^{*}(x))]_{i}\succeq 0$. Moreover, since $\bar{\lambda}^{*}(x)>0$, we have $\nabla^{2}_{yy}g(x,y^{*}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\succ 0$. Additionally, $\nabla_{y}\bar{h}(x,y^{*}(x))$ has full row rank. Thus, according to Lemma A.2 in [20], $H$ is invertible. We can calculate $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ with the following equation.
$\begin{bmatrix}\nabla y^{*}(x)\\ \nabla\bar{\lambda}^{*}(x)\end{bmatrix}=H^{-1}\begin{bmatrix}-\nabla^{2}_{xy}g(x,y^{*}(x))-\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ -\nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}$ (20)
According to [9], we have
$\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x)).$
Theorem˜4.9
When ˜3.1, 3.2, 3.3, 4.6 hold, and $\delta\leq\mu_{g}/(2l_{f,1})$, if ˜4.8 holds for a given $x$, we have
$\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|\leq O(\delta).$
Proof.
With $\lambda\in\mathbb{R}^{d_{h}}$, we have the following Lagrangian function
$\mathcal{L}_{\delta}(x,y,\lambda)=g_{\delta}(x,y)+\lambda^{\top}h(x,y).$
Denote $\lambda_{\delta}^{*}(x)$ as the optimal Lagrangian multiplier. We have the following KKT conditions.
$\nabla_{y}g_{\delta}(x,y^{*}_{\delta}(x))+\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)=0$
$\bar{h}(x,y^{*}_{\delta}(x))=0.$
Differentiating the KKT conditions with respect to $\delta$, we have
$\nabla_{y}f(x,y^{*}_{\delta}(x))+\nabla^{2}_{yy}g_{\delta}(x,y^{*}_{\delta}(x))\frac{d}{d\delta}y^{*}_{\delta}(x)+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)\frac{d}{d\delta}y^{*}_{\delta}(x)$
$+\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)=0$
$\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))\frac{d}{d\delta}y^{*}_{\delta}(x)=0$
We have the following equation:
$\begin{bmatrix}\nabla^{2}_{yy}g_{\delta}(x,y^{*}_{\delta}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)&\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\\ \nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))&0\end{bmatrix}\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}$
$=$ $\begin{bmatrix}-\nabla_{y}f(x,y^{*}_{\delta}(x))\\ 0\end{bmatrix}$. (21)
Denote
$H_{\delta}=\begin{bmatrix}\nabla^{2}_{yy}g_{\delta}(x,y^{*}_{\delta}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)&\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\\ \nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))&0\end{bmatrix}$. (22)
We can notice that $H_{\delta=0}=H$, where $H$ is defined in (19).
Then, we have
$\lim_{\delta\to 0}\frac{[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)]-[\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)]}{\delta}$
$=$ $\begin{bmatrix}\nabla^{2}_{xy}g(x,y^{*}(x))+\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ \nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}^{\top}\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}\Bigg|_{\delta=0}$
$=$ $\begin{bmatrix}\nabla^{2}_{xy}g(x,y^{*}(x))+\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ \nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}^{\top}H^{-1}\begin{bmatrix}-\nabla_{y}f(x,y^{*}(x))\\ 0\end{bmatrix}$
$=$ $T^{\top}H^{-1}\begin{bmatrix}-\nabla_{y}f(x,y^{*}(x))\\ 0\end{bmatrix}$
where we denote
$T=\begin{bmatrix}\nabla^{2}_{xy}g(x,y^{*}(x))+\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ \nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}$.
Note that according to (20), we have
$\nabla y^{*}(x)=-\begin{bmatrix}1&0\end{bmatrix}H^{-1}T$
$(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))=T^{\top}H^{-1}\begin{bmatrix}-\nabla_{y}f(x,y^{*}(x))\\ 0\end{bmatrix}$
Thus,
$\lim_{\delta\to 0}\frac{[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)]-[\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)]}{\delta}$
$=$ $(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (23)
Note that, according to Lemma˜4.4, we have
$\nabla\Phi_{\delta}(x)=$ $\frac{[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)]-[\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)]}{\delta}$
$+\nabla_{x}f(x,y^{*}_{\delta}(x)).$
Then, we consider to bound $\nabla\Phi_{\delta}(x)-\nabla\Phi(x)$.
$\nabla\Phi_{\delta}(x)-\nabla\Phi(x)$
$=$ $\nabla_{x}f(x,y^{*}_{\delta}(x))-\nabla_{x}f(x,y^{*}(x))$
$+\frac{[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)]-[\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)]}{\delta}$
$-(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$
$=$ $\nabla_{x}f(x,y^{*}_{\delta}(x))-\nabla_{x}f(x,y^{*}(x))$
$+\frac{[\nabla_{x}g(x,y^{*}_{\delta}(x))+\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)]-[\nabla_{x}g(x,y^{*}(x))+\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)]}{\delta}$
$-[\nabla^{2}_{xy}g(x,y^{*}(x))+\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)]\frac{d}{d\delta}y^{*}_{\delta}(x)|_{\delta=0}-\nabla_{x}\bar{h}(x,y^{*}(x))\frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)|_{\delta=0}$ (24)
where the last equality is due to (23).
For the first term in (24), we have
$\|\nabla_{x}f(x,y^{*}_{\delta}(x))-\nabla_{x}f(x,y^{*}(x))\|\leq l_{f,1}\|y^{*}_{\delta}(x)-y^{*}(x)\|\leq\frac{2\delta l_{f,0}l_{f,1}}{\mu_{g}}$
For the remaindering terms in (24), we have
$\frac{\nabla_{x}g(x,y^{*}_{\delta}(x))-\nabla_{x}g(x,y^{*}(x))}{\delta}-\nabla^{2}_{xy}g(x,y^{*}(x))\frac{d}{d\delta}y^{*}_{\delta}(x)|_{\delta=0}$
$+\frac{\nabla_{x}h(x,y^{*}_{\delta}(x))\lambda^{*}_{\delta}(x)-\nabla_{x}h(x,y^{*}(x))\lambda^{*}(x)}{\delta}-\nabla^{2}_{xy}h(x,y^{*}(x))\frac{d}{d\delta}\lambda^{*}_{\delta}(x)|_{\delta=0}$
$-\nabla_{x}\bar{h}(x,y^{*}(x))\frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)|_{\delta=0}$
$=$ $\frac{1}{\delta}\int_{t=0}^{\delta}\left(\nabla^{2}_{xy}g(x,y^{*}_{t}(x))\frac{d}{dt}y^{*}_{t}(x)-\nabla^{2}_{xy}g(x,y^{*}(x))\frac{d}{d\delta}y^{*}_{\delta}(x)|_{\delta=0}\right)dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\nabla^{2}_{xy}h(x,y^{*}_{t}(x))\frac{d}{dt}y^{*}_{t}(x)+\nabla_{x}h(x,y^{*}_{t}(x))\frac{d}{dt}\lambda^{*}_{t}(x)dt$
$-\frac{1}{\delta}\int_{t=0}^{\delta}\left(\nabla^{2}_{xy}h(x,y^{*}(x))\frac{d}{ds}y^{*}_{s}(x)+\nabla_{x}h(x,y^{*}(x))\frac{d}{ds}\lambda^{*}_{s}(x)\right)|_{s=0}dt$
$=$ $\frac{1}{\delta}\int_{t=0}^{\delta}\left[\nabla^{2}_{xy}g(x,y^{*}_{t}(x))-\nabla^{2}_{xy}g(x,y^{*}(x))\right]\frac{d}{dt}y^{*}_{t}(x)dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\nabla^{2}_{xy}g(x,y^{*}(x))\left[\frac{d}{dt}y^{*}_{t}(x)-\frac{d}{ds}y^{*}_{s}(x)|_{s=0}\right]dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\left[\nabla^{2}_{xy}h(x,y^{*}_{t}(x))-\nabla^{2}_{xy}h(x,y^{*}(x))\right]\frac{d}{dt}y^{*}_{t}(x)dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\nabla^{2}_{xy}h(x,y^{*}(x))\left[\frac{d}{dt}y^{*}_{t}(x)-\frac{d}{ds}y^{*}_{s}(x)|_{s=0}\right]dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\left[\nabla_{x}h(x,y^{*}_{t}(x))-\nabla_{x}h(x,y^{*}(x))\right]\frac{d}{dt}\lambda^{*}_{t}(x)dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\nabla_{x}h(x,y^{*}(x))\left[\frac{d}{dt}\lambda^{*}_{t}(x)-\lambda^{*}_{s}(x)|_{s=0}\right]dt$ (25)
For the first term in (25), we have
$\left\|\frac{1}{\delta}\int_{t=0}^{\delta}[\nabla^{2}_{xy}g(x,y^{*}_{t}(x))-\nabla^{2}_{xy}g(x,y^{*}(x))]\frac{d}{dt}y^{*}_{t}(x)dt\right\|$
$\leq$ $\frac{1}{\delta}\int_{t=0}^{\delta}\|\frac{d}{dt}y^{*}_{t}(x)\|\cdot l_{g,2}\|y_{t}^{*}(x)-y^{*}(x)\|\cdot dt$
$\leq$ $\frac{1}{\delta}\int_{t=0}^{\delta}l_{g,2}C_{y}^{2}\cdot t\cdot dt$
$=$ $\delta l_{g,2}C_{y}^{2}/2$
where the last equality is due to Lemma˜C.1.
Similarly, for the third, fifth terms in (25), we have
$\Bigg\|\frac{1}{\delta}\int_{t=0}^{\delta}\left[\nabla^{2}_{xy}h(x,y^{*}_{t}(x))-\nabla^{2}_{xy}h(x,y^{*}(x))\right]\frac{d}{dt}y^{*}_{t}(x)dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\left[\nabla_{x}h(x,y^{*}_{t}(x))-\nabla_{x}h(x,y^{*}(x))\right]\frac{d}{dt}\lambda^{*}_{t}(x)dt\Bigg\|$
$\leq$ $\delta(l_{h,2}+l_{h,1})C_{y}^{2}/2$
For the second term in (25), we have
$\|\frac{1}{\delta}\int_{t=0}^{\delta}\nabla^{2}_{xy}(x,y^{*}(x)\cdot\left(\frac{d}{dt}y^{*}_{t}(x)-\frac{d}{ds}y^{*}_{s}(x)|_{s=0}\right)dt\|$
$\leq$ $\frac{1}{\delta}\|\nabla^{2}_{xy}(x,y^{*}(x)\|\cdot\int_{t=0}^{\delta}\int_{s=0}^{t}\|\frac{d^{2}}{ds^{2}}y^{*}_{s}(x)\|dsdt$
$\leq$ $\frac{1}{\delta}\|\nabla^{2}_{xy}(x,y^{*}(x)\|\cdot\max_{s\in[0,\delta]}\|\frac{d^{2}}{ds^{2}}y^{*}_{s}(x)\|\cdot\delta^{2}$
$\leq$ $\delta l_{g,1}L_{y}$
where the last equality is due to Lemma˜C.1.
Similarly, for the fourth, sixth terms in (25), we have
$\Bigg\|\frac{1}{\delta}\int_{t=0}^{\delta}\nabla^{2}_{xy}g(x,y^{*}(x))\left[\frac{d}{dt}y^{*}_{t}(x)-\frac{d}{ds}y^{*}_{s}(x)|_{s=0}\right]dt$
$+\frac{1}{\delta}\int_{t=0}^{\delta}\nabla_{x}h(x,y^{*}(x))\left[\frac{d}{dt}\lambda^{*}_{t}(x)-\lambda^{*}_{s}(x)|_{s=0}\right]dt\Bigg\|$
$\leq$ $\delta(l_{h,1}+l_{h,0})L_{y}$
Therefore,
$\|\nabla\Phi_{\delta}(x)-\nabla\Phi(x)\|\leq O(\delta).$
Lemma C.1.
When ˜3.1, 3.2, 3.3, 4.6 hold, and $\delta\leq\mu_{g}/(2l_{f,1})$, if ˜4.8 holds for a given $x$, we have
$\left\|\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}\right\|\leq C_{y}$
$\left\|\begin{bmatrix}\frac{d^{2}}{d\delta^{2}}y^{*}_{\delta}(x)\\ \frac{d^{2}}{d\delta^{2}}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}\right\|\leq L_{y}$
where $C_{y},L_{y}$ are $O(1)$ constants.
Proof.
According to (21), we have
$\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}=H_{\delta}^{-1}p_{\delta},$
where
$H_{\delta}=\begin{bmatrix}\nabla^{2}_{yy}g_{\delta}(x,y^{*}_{\delta}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)&\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\\ \nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))&0\end{bmatrix}$,
$p_{\delta}=\begin{bmatrix}-\nabla_{y}f(x,y^{*}_{\delta}(x))\\ 0\end{bmatrix}$.
According to Lemma˜C.2, we have
$\left\|\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta}(x)\end{bmatrix}\right\|\leq\left\|H_{\delta}^{-1}p_{\delta}\right\|\leq C_{H}l_{f,0}=C_{y}.$
Denote
$\left\|H_{\delta_{2}}-H_{\delta_{1}}\right\|=\left\|\begin{bmatrix}B&D^{T}\\ D&0\end{bmatrix}\right\|$
where
$B=$ $\nabla^{2}_{yy}g_{\delta_{2}}(x,y^{*}_{\delta_{2}}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{2}}(x))^{\top}\bar{\lambda}^{*}_{\delta_{2}}(x)-\nabla^{2}_{yy}g_{\delta_{1}}(x,y^{*}_{\delta_{1}}(x))-\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{1}}(x))^{\top}\bar{\lambda}^{*}_{\delta_{1}}(x)$
$=$ $\delta_{2}\nabla^{2}_{yy}f(x,y^{*}_{\delta_{2}}(x))-\delta_{1}\nabla^{2}_{yy}f(x,y^{*}_{\delta_{1}}(x))+\nabla^{2}_{yy}g(x,y^{*}_{\delta_{2}}(x))-\nabla^{2}_{yy}g(x,y^{*}_{\delta_{1}}(x))$
$+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{2}}(x))^{\top}\bar{\lambda}^{*}_{\delta_{2}}(x)-\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{1}}(x))^{\top}\bar{\lambda}^{*}_{\delta_{1}}(x)$
$=$ $\delta_{1}(\nabla^{2}_{yy}f(x,y^{*}_{\delta_{2}}(x))-\nabla^{2}_{yy}f(x,y^{*}_{\delta_{1}}(x)))+\nabla^{2}_{yy}f(x,y^{*}_{\delta_{2}}(x))(\delta_{2}-\delta_{1})+\nabla^{2}_{yy}g(x,y^{*}_{\delta_{2}}(x))$
$-\nabla^{2}_{yy}g(x,y^{*}_{\delta_{1}}(x))+[\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{2}}(x))-\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{1}}(x))]^{\top}\bar{\lambda}^{*}_{\delta_{2}}(x)$
$+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta_{1}}(x))^{\top}[\bar{\lambda}^{*}_{\delta_{2}}(x)-\bar{\lambda}^{*}_{\delta_{1}}(x)].$
Thus
$\|B\|\leq\left(\frac{2\mu_{g}l_{f,2}}{l_{f,1}}C_{y}+l_{f,1}+l_{g,2}C_{y}+\Lambda l_{h,2}C_{y}+l_{f,1}C_{y}\right)|\delta_{2}-\delta_{1}|$
$\|D\|=\|\nabla_{y}\bar{h}(x,y^{*}_{\delta_{2}}(x))-\nabla_{y}\bar{h}(x,y^{*}_{\delta_{1}}(x))\|\leq l_{h,1}(y^{*}_{\delta_{2}}(x)-y^{*}_{\delta_{1}}(x))\leq l_{h,1}C_{y}|\delta_{1}-\delta_{2}|$
Therefore, we have
$\left\|H_{\delta_{2}}-H_{\delta_{1}}\right\|\leq M_{H}|\delta_{1}-\delta_{2}|$
Moreover, we have
$\left\|\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta_{1}}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta_{1}}(x)\end{bmatrix}-\begin{bmatrix}\frac{d}{d\delta}y^{*}_{\delta_{2}}(x)\\ \frac{d}{d\delta}\bar{\lambda}^{*}_{\delta_{2}}(x)\end{bmatrix}\right\|$
$=$ $\|H_{\delta_{1}}^{-1}p_{\delta_{1}}-H_{\delta_{2}}^{-1}p_{\delta_{2}}\|$
$=$ $\|(H_{\delta_{1}}^{-1}-H_{\delta_{2}}^{-1})p_{\delta_{1}}+H_{\delta_{2}}^{-1}(p_{\delta_{1}}-p_{\delta_{2}})\|$
$\leq$ $\|(H_{\delta_{1}}^{-1}(H_{\delta_{2}}-H_{\delta_{1}})H_{\delta_{2}}^{-1})\|\|p_{\delta_{1}}\|+\|H_{\delta_{2}}^{-1}\|\|p_{\delta_{1}}-p_{\delta_{2}}\|$
$\leq$ $\left(C_{H}^{2}l_{f,0}M_{H}^{2}+C_{H}l_{f,1}\right)|\delta_{1}-\delta_{2}|=L_{y}|\delta_{1}-\delta_{2}|$
where $L_{y}=\left(C_{H}^{2}l_{f,0}M_{H}^{2}+C_{H}l_{f,1}\right)$.
Lemma C.2.
For $H_{\delta}$ defined in (22), we have
$\|H_{\delta}^{-1}\|\leq C_{H},$
where $C_{H}$ is an $O(1)$ constant depending on $\mu_{g},l_{g,1},l_{f,1},d_{h},l_{h,1},s_{\min},s_{\max},\Lambda$.
Proof.
Denote $A=\nabla^{2}_{yy}g_{\delta}(x,y^{*}_{\delta}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}_{\delta}(x))^{\top}\bar{\lambda}^{*}_{\delta}(x)$, $C=\nabla_{y}\bar{h}(x,y^{*}_{\delta}(x))$. We have
$H_{\delta_{1}}^{-1}=\begin{bmatrix}A^{-1}+A^{-1}C^{\top}(CA^{-1}C^{\top})^{-1}CA^{-1}&-A^{-1}C^{\top}(CA^{-1}C^{\top})^{-1}\\ -(CA^{-1}C^{\top})^{-1}CA^{-1}&(CA^{-1}C^{\top})^{-1}\end{bmatrix}$.
According to ˜4.8, we know that $\|\lambda^{*}_{\delta}(x)\|\leq\Lambda$. Thus, $0\leq[\lambda^{*}_{\delta}(x)]_{i}\leq\Lambda$. We have
$\|A^{-1}\|\leq\frac{2}{\mu_{g}}$
$\|A\|\leq L_{g}+d_{h}l_{h,1}\Lambda$
$\|C\|\leq s_{\max}$
$\|C^{-1}\|\leq\frac{1}{s_{\min}}$
Denote
$H_{\delta_{1}}^{-1}=\begin{bmatrix}B&D^{T}\\ D&E\end{bmatrix}$.
we have $\|B\|\leq C_{B},\|D\|\leq C_{D},\|E\|\leq C_{E}$ and so that $\|H_{\delta_{1}}^{-1}\|\leq C_{H}$, where $C_{B},C_{D},C_{E},C_{H}$ are $O(1)$ constants depending on $\mu_{g},l_{g,1},l_{f,1},d_{h},l_{h,1},s_{\min},s_{\max},\Lambda$.
Appendix D Proofs of Theorem˜5.4 and Corollary˜5.5
In this section, we provide proofs for Theorem˜5.4 and Corollary˜5.5. We first introduce the additional notations and lemmas that will be used in this section.
Notations
$K(x,y^{\prime},z^{\prime},u,v)=\phi_{\delta}(x,y,z)+u^{\top}(A^{\prime}y^{\prime}-b)-v^{\top}(A^{\prime}z^{\prime}-b)+\frac{\rho_{1}}{2}\|A^{\prime}y^{\prime}-b\|^{2}-\frac{\rho_{2}}{2}\|A^{\prime}z^{\prime}-b\|^{2}$
$L_{K}=l_{\delta}+2l_{g,1}+\max\{\rho_{1},\rho_{2}\}\sigma_{\max}^{2}(A^{\prime})$
$\mu_{y}=\min\{\mu_{g}-l_{\delta}-\rho_{1}\sigma_{\max}^{2}(A),\frac{\rho_{1}}{2}\}$
$\mu_{z}=\min\{\mu_{g}-\rho_{2}\sigma_{\max}^{2}(A),\frac{\rho_{2}}{2}\}$
$y^{*}_{\delta}(x)=\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)$
$z^{*}(x)=\mathop{\rm argmin}_{z\in\mathcal{Y}(x)}g(x,z)$
$y^{\prime*}_{\delta}(x,u)=\mathop{\rm argmin}_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v)$
$z^{\prime*}(x,v)=\mathop{\rm argmax}_{z^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v)$
$[y^{*}_{\delta}(x,u)^{\top},\alpha^{*}_{\delta}(x,u)^{\top}]^{\top}=y^{\prime*}_{\delta}(x,u)$
$[z^{*}(x,v)^{\top},\beta^{*}(x,v)^{\top}]^{\top}=z^{\prime*}(x,v)$
$d(x,z^{\prime},u,v)=K(x,y^{\prime*}_{\delta}(x,u),z^{\prime},u,v)$
$q(x,v)=\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x,v))-v^{\top}(A^{\prime}z^{\prime*}(x,v)-b)-\frac{\rho_{2}}{2}\|A^{\prime}z^{\prime*}(x,v)-b\|^{2}$
$V_{t}=\frac{1}{4}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{t},v_{t})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
Lemma D.1.
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$ and $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $K(x,y^{\prime},z^{\prime},u,v)$ is $\mu_{y}$-strongly convex w.r.t. $y^{\prime}$, $\mu_{z}$-strongly concave w.r.t. $z^{\prime}$, and $L_{K}$-smooth w.r.t. $x,y^{\prime},z^{\prime}$.
Proof.
According to Lemma˜A.4, we know that $K(x,y^{\prime},z^{\prime},u,v)$ is $\mu_{y}$-strongly convex w.r.t. $y^{\prime}$, $\mu_{z}$-strongly concave w.r.t. $z^{\prime}$. Moreover
$\nabla_{x}K(x,y^{\prime},z^{\prime},u,v)=\nabla_{x}\phi_{\delta}(x,y,z)$
$\nabla_{y}K(x,y^{\prime},z^{\prime},u,v)=\nabla_{y}\phi_{\delta}(x,y,z)+A^{\top}u+\rho_{1}A^{\top}(A^{\prime}y^{\prime}-b)$
$\nabla_{z}K(x,y^{\prime},z^{\prime},u,v)=\nabla_{z}\phi_{\delta}(x,y,z)-A^{\top}v-\rho_{2}A^{\top}(A^{\prime}z^{\prime}-b)$
Thus, $K(x,y^{\prime},z^{\prime},u,v)$ is $L_{K}$-smooth w.r.t. $x,y^{\prime},z^{\prime}$. ∎
Lemma D.2.
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$ and $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y},\eta_{z}\leq 1/L_{K}$, we have
$\|y^{\prime*}_{\delta}(x,u_{1})-y^{\prime*}_{\delta}(x,u_{2})\|\leq\sigma_{yu}\|u_{1}-u_{2}\|$ (26)
$\|y^{\prime*}_{\delta}(x_{1},u)-y^{\prime*}_{\delta}(x_{2},u)\|\leq\sigma_{yx}\|x_{1}-x_{2}\|$ (27)
$\|z^{\prime*}(x,v_{1})-z^{\prime*}(x,v_{2})\|\leq\sigma_{zv}\|v_{1}-v_{2}\|$ (28)
$\|z^{\prime*}(x_{1},v)-z^{\prime*}(x_{2},v)\|\leq\sigma_{zx}\|x_{1}-x_{2}\|$ (29)
$\|y^{*}_{\delta}(x_{1})-y^{*}_{\delta}(x_{2})\|\leq\sigma_{ys}\|x_{1}-x_{2}\|$ (30)
$\|z^{*}(x_{1})-z^{*}(x_{2})\|\leq\sigma_{zs}\|x_{1}-x_{2}\|$ (31)
$\|y^{\prime*}_{\delta}(x,u)-y^{\prime*}_{\delta}(x)\|\leq\sigma_{y}\|A^{\prime}y^{\prime*}_{\delta}(x,u)-b\|$ (32)
$\|z^{\prime*}(x,v)-z^{\prime*}(x)\|\leq\sigma_{z}\|A^{\prime}z^{\prime*}(x,v)-b\|$ (33)
$\|y^{\prime*}_{\delta}(x,u)-y^{\prime}\|\leq\sigma_{ye}\|\nabla_{y}K(x,y^{\prime},z^{\prime},u,v)\|+\sigma_{\alpha}\|\alpha-\Pi_{-}(\alpha-\eta_{y}\nabla_{\alpha}K(x,y^{\prime},z^{\prime},u,v))\|$ (34)
$\|z^{\prime*}(x,v)-z^{\prime}\|\leq\sigma_{ze}\|\nabla_{z}K(x,y^{\prime},z^{\prime},u,v)\|+\sigma_{\beta}\|\beta-\Pi_{-}(\beta+\eta_{z}\nabla_{\beta}K(x,y^{\prime},z^{\prime},u,v))\|$ (35)
$\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|\leq\frac{2}{\mu_{y}\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|$ (36)
$\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|\leq\frac{2}{\mu_{z}\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|$ (37)
where $\sigma_{yu}=\frac{\sigma_{\max}(A^{\prime})}{\mu_{y}}$, $\sigma_{yx}=\frac{L_{K}}{\mu_{y}}$, $\sigma_{zv}=\frac{\sigma_{\max}(A^{\prime})}{\mu_{z}}$, $\sigma_{zx}=\frac{L_{K}}{\mu_{z}}$, $\sigma_{ye}=\frac{2}{\mu_{y}}$, $\sigma_{ze}=\frac{2}{\mu_{z}}$, $\sigma_{\alpha}=\frac{2}{\mu_{y}\eta_{y}}$, $\sigma_{\beta}=\frac{2}{\mu_{z}\eta_{z}}$, $\sigma_{ys}=\frac{2L_{K}}{\mu_{g}}$, $\sigma_{zs}=\frac{L_{K}}{\mu_{g}}$,
$\sigma_{y}=\frac{\sqrt{2}(\bar{\theta}L_{K}^{2}+1)}{\mu_{y}}$
$\sigma_{z}=\frac{\sqrt{2}(\bar{\theta}L_{K}^{2}+1)}{\mu_{z}}$
$M=\begin{pmatrix}A^{\prime\top}&G^{\top}\\ 0&I\end{pmatrix}$
$G=\begin{pmatrix}0_{d_{y}\times d_{h}}&0\\ 0&I_{d_{h}}\end{pmatrix}$
$\bar{\theta}=\max_{\bar{M}\in\mathcal{B}(M)}\sigma^{2}_{\max}(\bar{M})/\sigma^{4}_{\min}(\bar{M})$
$\bar{M}$ is the set of all submatrices of $M$ with full row rank.
Proof.
(26), (27), (28), (29), (30), (31) is due to Lemma˜A.2. (34), (35), (36), (37) is due to Lemma˜A.1. (32), (33) is due to Lemma˜A.3. ∎
D.1 Potential function
In this subsection, we will prove the following descent lemma for $V_{t}$.
Lemma D.3.
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$, $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y}=1/(4L_{K})$, $\eta_{z}=2/(L_{K}+4L_{d})$, $\eta_{x}=\min\{\eta_{y}\mu_{y}^{2}/(512L_{\phi}^{2}),\eta_{z}\mu_{z}^{2}/(96L_{\phi}^{2}),\eta_{u}/(64\sigma_{y}^{2}L_{\phi}^{2}),\eta_{v}/(4\sigma_{z}^{2}L_{\phi}^{2}),2/(L_{K}+4L_{d}+8L_{q})\}$, $\eta_{u}=\eta_{y}\mu_{y}^{2}/(32\sigma_{\max}^{2}(A))$, $\eta_{v}=\eta_{z}\mu_{z}^{2}/(32\sigma_{\max}^{2}(A))$, we have
$V_{t}-V_{t+1}\geq$ $\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{16\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\frac{1}{8\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}+\frac{\eta_{v}}{2}\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x_{t})\|^{2}$
Thus,
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}}(V_{0}-\min_{t}V_{t})$ (38)
Proof.
First, for function $d$, we have
$d(x_{t},z^{\prime}_{t},u_{t+1},v_{t+1})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
$=$ $K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t}),z^{\prime}_{t},u_{t},v_{t})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t},v_{t})$
$=$ $(u_{t+1}-u_{t})^{\top}(A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b)-(v_{t+1}-v_{t})^{\top}(A^{\prime}z^{\prime}_{t}-b)$
$=$ $-\eta_{v}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{u}(A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b)^{\top}(A^{\prime}y^{\prime}_{t}-b)$
Note that
$\nabla_{x}d(x,z^{\prime},u,v)=\nabla_{x}\phi_{\delta}(x,y^{*}_{\delta}(x,u),z)$
$\nabla_{z^{\prime}}d(x,z^{\prime},u,v)=\begin{bmatrix}\nabla_{z}g(x,z)-A^{\top}v-\rho_{2}A^{\top}(A^{\prime}z^{\prime}-b)\\ v+(A^{\prime}z^{\prime}-b)\end{bmatrix}$
Thus, according to Lemma˜D.2, we know that $\nabla_{x}d(x,z^{\prime},u,v)$ is $(L_{\phi}+L_{\phi}\sigma_{yx})$-continuous w.r.t. $x,z^{\prime}$ and $\nabla_{z}^{\prime}d(x,z^{\prime},u,v)$ is $L_{K}$-continuous w.r.t. $x,z^{\prime}$. Define $L_{d}=\max\{L_{\phi}+L_{\phi}\sigma_{yx},L_{K}\}$. We have
$d(x_{t+1},z^{\prime}_{t+1},u_{t+1},v_{t+1})-d(x_{t},z^{\prime}_{t},u_{t+1},v_{t+1})$
$\geq$ $\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),x_{t+1}-x_{t}\rangle$
$+\langle\nabla_{z^{\prime}}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),z^{\prime}_{t+1}-z^{\prime}_{t}\rangle$
$-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
$\geq$ $\langle\nabla_{x}\phi_{\delta}(x_{t},y_{\delta}^{*}(x_{t},u_{t+1}),z_{t}),x_{t+1}-x_{t}\rangle+\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
Then, for function $q$, we have
$q(x_{t},v_{t})-q(x_{t},v_{t+1})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t}),u_{t},v_{t})-K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t+1})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t})-K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t+1})$
$\geq$ $\eta_{v}(A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b)^{\top}(A^{\prime}z^{\prime}_{t}-b)$
Note that
$\nabla_{x}q(x,v)=\nabla_{x}\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x,v))$
Thus, according to Lemma˜D.2, $q(\cdot,v)$ is $L_{q}=(L_{\phi}+L_{\phi}\sigma_{zx}+L_{\phi}\sigma_{ys})$-smooth. We have
$q(x_{t},v_{t+1})-q(x_{t+1},v_{t+1})$
$\geq$ $\langle\nabla_{x}\phi_{\delta}(x_{t},y_{\delta}^{*}(x_{t}),z^{*}(x_{t},v_{t+1})),x_{t}-x_{t+1}\rangle-\frac{L_{q}}{2}(\|x_{t+1}-x_{t}\|^{2})$
Finally, for function $K$, we have
$K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})-K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})=-\eta_{u}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\eta_{v}\|A^{\prime}z^{\prime}_{t}-b\|^{2}$
and
$K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t+1},y^{\prime}_{t+1},z^{\prime}_{t+1},u_{t+1},v_{t+1})$
$\geq$ $\frac{1}{\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$-\frac{L_{K}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}).$
Thus, for $V_{t}$, we have
$V_{t}-V_{t+1}$
$\geq$ $\langle\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t},u_{t+1}),z_{t}),x_{t+1}-x_{t}\rangle+\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\eta_{u}(A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b)^{\top}(A^{\prime}y^{\prime}_{t}-b)-\eta_{v}\|A^{\prime}z^{\prime}_{t}-b\|^{2}-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
$+2\langle\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t}),z^{*}(x_{t},v_{t+1})),x_{t}-x_{t+1}\rangle$
$+2\eta_{v}(A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b)^{\top}(A^{\prime}z^{\prime}_{t}-b)-L_{q}(\|x_{t+1}-x_{t}\|^{2})$
$+\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{4\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\frac{1}{4\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}-\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}$
$-\frac{L_{K}}{8}(\|x_{t+1}-x_{t}\|^{2}+\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
$\geq$ $-\|\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t},u_{t+1}),z_{t})-\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t}),z^{*}(x_{t},v_{t+1}))\|\|x_{t+1}-x_{t}\|$
$+\left(\frac{\eta_{u}}{2}-\frac{\eta_{u}}{4}\right)\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-\frac{\eta_{u}}{2}\|A^{\prime}y^{\prime}_{t}-A^{\prime}y_{\delta}^{\prime*}(x_{t},u_{t+1})\|^{2}$
$+\left(\eta_{v}+\frac{\eta_{v}}{4}-\eta_{v}\right)\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\eta_{v}\|A^{\prime}z^{\prime}_{t}-A^{\prime}z^{\prime*}(x_{t},v_{t+1})\|^{2}$
$-\frac{\eta_{x}}{2}\|\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t}),z^{*}(x_{t},v_{t+1}))-\nabla_{x}\phi_{\delta}(x_{t},y_{t},z_{t})\|^{2}$
$+\frac{1}{2\eta_{x}}\|x_{t}-x_{t+1}\|^{2}+\frac{\eta_{x}}{2}\|\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t}),z^{*}(x_{t},v_{t+1}))\|^{2}$
$+\left(\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{\eta_{z}}-\frac{1}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}$
$\geq$ $-\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}-2\eta_{x}L_{\phi}^{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime*}_{\delta}(x_{t})\|^{2}-2\eta_{x}L_{\phi}^{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-\frac{\eta_{u}\sigma_{\max}^{2}(A)}{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}$
$+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\eta_{v}\sigma_{\max}^{2}(A)\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$-2\eta_{x}L_{\phi}^{2}\|y^{\prime*}_{\delta}(x_{t})-y^{\prime*}_{\delta}(x_{t},u_{t+1})\|^{2}-2\eta_{x}L_{\phi}^{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}-\eta_{x}L_{\phi}^{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{x}}{4}\|\nabla_{x}\phi_{\delta}(x_{t},y^{*}_{\delta}(x_{t}),z^{*}(x_{t}))\|^{2}-\frac{\eta_{x}L_{\phi}^{2}}{2}\|z^{\prime*}(x_{t})-z^{\prime*}(x_{t},v_{t+1})\|^{2}+\frac{1}{2\eta_{x}}\|x_{t}-x_{t+1}\|^{2}$
$+\left(\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2},$
and
$V_{t}-V_{t+1}$
$\geq$ $\left(\frac{1}{2\eta_{x}}+\frac{1}{4\eta_{x}}-\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\left(4\eta_{x}L_{\phi}^{2}+\frac{\eta_{u}\sigma_{\max}^{2}(A)}{2}\right)\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}-\left(3\eta_{x}L_{\phi}^{2}+\eta_{v}\sigma_{\max}^{2}(A)\right)\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-4\eta_{x}L_{\phi}^{2}\|y_{\delta}^{\prime*}(x_{t},u_{t+1})-y^{\prime*}_{\delta}(x_{t})\|^{2}$
$+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\frac{\eta_{x}L_{\phi}^{2}}{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime*}(x_{t})\|^{2}$
$+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x_{t})\|^{2}$
$\geq$ $\left(\frac{1}{2\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}-\frac{16\eta_{x}L_{\phi}^{2}+2\eta_{u}\sigma_{\max}^{2}(A)}{\mu_{y}^{2}\eta_{y}^{2}}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-\frac{12\eta_{x}L_{\phi}^{2}+4\eta_{v}\sigma_{\max}^{2}(A)}{\mu_{z}^{2}\eta_{z}^{2}}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{\eta_{u}}{2}-4\sigma_{y}^{2}\eta_{x}L_{\phi}^{2}\right)\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}$
$+\left(\eta_{v}-\frac{\sigma_{z}^{2}\eta_{x}L_{\phi}^{2}}{2}\right)\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x_{t})\|^{2}$
where the last equality is due to Lemma˜D.2.
Thus, when $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$, $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y}=1/(4L_{K})$, $\eta_{z}=2/(L_{K}+4L_{d})$, $\eta_{x}=\min\{\eta_{y}\mu_{y}^{2}/(512L_{\phi}^{2}),\eta_{z}\mu_{z}^{2}/(96L_{\phi}^{2}),\eta_{u}/(64\sigma_{y}^{2}L_{\phi}^{2}),\eta_{v}/(4\sigma_{z}^{2}L_{\phi}^{2}),2/(L_{K}+4L_{d}+8L_{q})\}$, $\eta_{u}=\eta_{y}\mu_{y}^{2}/(32\sigma_{\max}^{2}(A))$, $\eta_{v}=\eta_{z}\mu_{z}^{2}/(32\sigma_{\max}^{2}(A))$, we have
$V_{t}-V_{t+1}$
$\geq$ $\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{16\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\frac{1}{8\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}+\frac{\eta_{v}}{2}\|A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x_{t})\|^{2}$ (39)
Thus,
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}}(V_{0}-\min_{t}V_{t})$ (40)
Proof of Theorem˜5.4
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$, $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y}=1/(4L_{K})$, $\eta_{z}=2/(L_{K}+4L_{d})$, $\eta_{x}=\min\{\eta_{y}\mu_{y}^{2}/(512L_{\phi}^{2}),\eta_{z}\mu_{z}^{2}/(96L_{\phi}^{2}),\eta_{u}/(64\sigma_{y}^{2}L_{\phi}^{2}),\eta_{v}/(4\sigma_{z}^{2}L_{\phi}^{2}),2/(L_{K}+4L_{d}+8L_{q})\}$, $\eta_{u}=\eta_{y}\mu_{y}^{2}/(32\sigma_{\max}^{2}(A))$, $\eta_{v}=\eta_{z}\mu_{z}^{2}/(32\sigma_{\max}^{2}(A))$, according to (40), we have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}}(V_{0}-\min_{t}V_{t})$ (41)
Note that
$V_{t}=$ $\frac{1}{4}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{t},v_{t})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
$\geq$ $2q(x_{t},v_{t})-\frac{3}{4}d(x_{t},z^{\prime}_{t},u_{t},v_{t})\geq\frac{5}{4}q(x_{t},v_{t})\geq\frac{5}{4}\phi_{\delta}(x_{t})\geq\frac{5\delta\Phi^{*}}{4}-\frac{5\delta^{2}l^{2}_{f,0}}{8\mu_{g}}$
Therefore, when $\delta=\Theta(\epsilon)$, with $T=O(\epsilon^{-4})$, we have $t\in[T]$, such that $\|\nabla\Phi_{\delta}(x_{t})\|=\|\frac{1}{\delta}\nabla\phi_{\delta}(x_{t})\|\leq\epsilon$.
Moreover, if we have $x_{0},y_{0},z_{0},u_{0},v_{0}$ such that
$\|y_{0}-y^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|\leq O(\delta)$
$\|A^{\prime}y_{0}^{\prime}-b\|\leq O(\delta)$
$\|z_{0}-z^{*}(x_{0})\|\leq O(\delta)$
$\|A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|\leq O(\delta)$
$\|A^{\prime}z_{0}^{\prime}-b\|\leq O(\delta)$
Then, we have
$V_{0}$
$=$ $\frac{1}{4}\bigg[\delta f(x_{0},y_{0})+(g(x_{0},y_{0})-g(x_{0},z_{0}))+u_{0}^{\top}(A^{\prime}y_{0}^{\prime}-b)-v_{0}^{\top}(A^{\prime}z_{0}^{\prime}-b)$
$+\frac{\rho_{1}}{2}\|A^{\prime}y_{0}^{\prime}-b\|^{2}-\frac{\rho_{2}}{2}\|A^{\prime}z_{0}^{\prime}-b\|^{2}\bigg]$
$+2\bigg[\delta f(x_{0},y^{*}_{\delta}(x_{0}))+(g(x_{0},y^{*}_{\delta}(x_{0}))-g(x_{0},z^{*}(x_{0},v_{0})))-v_{0}^{\top}(A^{\prime}z^{\prime*}(x_{0},v_{0})-b)$
$-\frac{\rho_{2}}{2}\|A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|^{2}\bigg]$
$-\bigg[\delta f(x_{0},y^{*}_{\delta}(x_{0},u_{0}))+(g(x_{0},y^{*}_{\delta}(x_{0},u_{0}))-g(x_{0},z_{0}))+u_{0}^{\top}(A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b)$
$-v_{0}^{\top}(A^{\prime}z_{0}^{\prime}-b)+\frac{\rho_{1}}{2}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|^{2}-\frac{\rho_{2}}{2}\|A^{\prime}z_{0}^{\prime}-b\|^{2}\bigg]$
$\leq$ $\frac{5\delta\Phi(x_{0})}{4}+O(1)l_{f,1}(\|y_{0}-y^{*}(x_{0})\|+\|y^{*}_{\delta}(x_{0})-y^{*}(x_{0})\|+\|y^{*}_{\delta}(x_{0},u_{0})-y^{*}(x_{0})\|)$
$+O(1)C_{g}(\|z_{0}-y_{0}\|+\|y^{*}_{\delta}(x_{0})-z^{*}(x_{0},v_{0})\|+\|y^{*}_{\delta}(x_{0},u_{0})-z_{0}\|)$
$+O(\|A^{\prime}z_{0}^{\prime}-b\|+\|A^{\prime}z_{0}^{\prime}-b\|^{2}+\|A^{\prime}y_{0}^{\prime}-b\|+\|A^{\prime}y_{0}^{\prime}-b\|^{2}$
$+\|A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|+\|A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|^{2}+\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|+\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|^{2})$
$\leq$ $\frac{5\delta\Phi(x_{0})}{4}+O(\delta)$
where $G=\max\{g(x_{0},y_{0}),g(x_{0},z_{0}),g(x_{0},y^{*}_{\delta}(x_{0}),g(x_{0},z^{*}(x_{0},v_{0})),g(x_{0},y^{*}_{\delta}(x_{0},u_{0})\}$, $\mathcal{C}=\{y\in\mathbb{R}^{d_{y}}|g(x_{0},y)\leq G\}$, $C_{g}=\sup_{y\in\mathcal{C}}\nabla_{y}g(x_{0},y)$. Since $g(x_{0},y)$ is strongly convex w.r.t $y$, its sub-level set $\mathcal{C}$ is compact and convex. Moreover, since $g$ is Lipschitz smoothness, its gradient in this compact set $\mathcal{C}$ is upper bounded by an $O(1)$ constant $C_{g}$.
We can notice that
$V_{0}-\min_{t}V_{t}\leq\frac{5\delta[\Phi(x_{0})-\Phi^{*}]}{4}+O(\delta)=O(\epsilon)$
and
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\Phi_{\delta}(x_{t})\|^{2}=\frac{1}{T\delta^{2}}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}\delta^{2}}(V_{0}-\min_{t}V_{t})=O\left(\frac{\epsilon^{-1}}{T}\right).$
Therefore, we can find an $\epsilon$-stationary point of $\Phi_{\delta}(x)$ with a complexity of $O(\epsilon^{-3})$.
Proof of Corollary˜5.5
For fixed $x_{0}$, define $W_{t}=\frac{1}{4}K(x_{0},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{0},v_{t})-d(x_{0},z^{\prime}_{t},u_{t},v_{t})$. According to (39), with appropriate parameters, we have
$W_{t}-W_{t+1}$
$\geq$ $\frac{1}{16\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\frac{1}{8\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|^{2}+\frac{\eta_{v}}{2}\|A^{\prime}z^{\prime*}(x_{0},v_{t+1})-b\|^{2}$
$+\frac{\eta_{u}}{4}\|A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|A^{\prime}z^{\prime}_{t}-b\|^{2}$
Thus, when $T=O(\epsilon^{2})$, we can find $t\in T$ such that
$\|y^{\prime}_{t+1}-y^{\prime}_{t}\|\leq\delta$
$\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|\leq\delta$
$\|A^{\prime}y^{\prime}_{t}-b\|\leq\delta$
$\|z^{\prime}_{t+1}-z^{\prime}_{t}\|\leq\delta$
$\|A^{\prime}z^{\prime*}(x_{0},v_{t+1})-b\|\leq\delta$
$\|A^{\prime}z^{\prime}_{t}-b\|\leq\delta$
Denote the active set at $y^{*}_{\delta}(x_{0})$ as $\mathcal{I}^{\alpha}$, $\mathcal{J}^{\alpha}=[d_{h}]/\mathcal{I}$. Define $\Delta^{\alpha}=\min_{i\in\mathcal{J}^{\alpha}}|[\alpha^{*}_{\delta}(x_{0})]_{i}|$. Denote the active set of $z^{*}(x_{0})$ as $\mathcal{I}^{\beta}$, $\mathcal{J}^{\beta}=[d_{h}]/\mathcal{I}$. Define $\Delta^{\beta}=\min_{i\in\mathcal{J}^{\beta}}|[\beta^{*}(x_{0})]_{i}|$. Set $\epsilon\leq\min\{\Delta^{\alpha}/(6\sigma_{\alpha}),\Delta^{\alpha}/(6\sigma_{y}),\Delta^{\beta}/(6\sigma_{\beta}),\Delta^{\beta}/(6\sigma_{z})\}$. According to Lemma˜D.4, we have
$\|u_{t+1}-u^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|v_{t+1}-v^{*}(x_{0})\|\leq O(\delta)$
Then, if we set $y^{\prime}_{0}=y^{\prime}_{t}$, $z^{\prime}_{0}=z^{\prime}_{t}$, $u_{0}=u_{t+1}$, $v_{0}=v_{t+1}$, with $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points, according to Theorem˜5.4, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$.
Thus, the total complexity is $O(\epsilon^{-3}+\epsilon^{-2})=O(\epsilon^{-3})$.
Lemma D.4.
For a fixed $x_{0}$, denote the active set at $y^{*}_{\delta}(x_{0})$ as $\mathcal{I}^{\alpha}$, $\mathcal{J}^{\alpha}=[d_{h}]/\mathcal{I}$. We have $\alpha^{*}_{\delta}(x_{0})]_{\mathcal{I}^{\alpha}}=0$ and $\alpha^{*}_{\delta}(x_{0})]_{\mathcal{J}^{\alpha}}<0$. Suppose $s^{\alpha}_{\min}=\sigma_{\min}(A_{\mathcal{I}^{\alpha}})>0$. Define $\Delta^{\alpha}=\min_{i\in\mathcal{J}^{\alpha}}|[\alpha^{*}_{\delta}(x_{0})]_{i}|$. When $\|y^{\prime}_{t+1}-y^{\prime}_{t}\|\leq\Delta^{\alpha}/(6\sigma_{\alpha}),\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|\leq\Delta^{\alpha}/(6\sigma_{y})$, we have
$\|u_{t+1}-u^{*}_{\delta}(x_{0})\|\leq$ $\sigma_{uyx0}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{u2x0}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|+\sigma_{u1x0}\|A^{\prime}y^{\prime}_{t}-b\|$
where $\sigma_{uyx0}=\frac{1}{\eta_{y}}+\frac{1+\sigma_{\max}}{\eta_{y}s_{\min}^{\alpha}}+\frac{L_{g}\sigma_{\alpha}}{s_{\min}^{\alpha}}$, $\sigma_{u2x0}=\frac{L_{g}\sigma_{y}}{s_{\min}^{\alpha}}$, $\sigma_{u1x0}=\frac{\rho_{1}\sigma_{\max}(A)}{s_{\min}^{\alpha}}$.
Similarly, for a fixed $x_{0}$, denote the active set of $z^{*}(x_{0})$ as $\mathcal{I}^{\beta}$, $\mathcal{J}^{\beta}=[d_{h}]/\mathcal{I}$. Suppose $s^{\beta}_{\min}=\sigma_{\min}(A_{\mathcal{I}^{\beta}})>0$. Define $\Delta^{\beta}=\min_{i\in\mathcal{J}^{\beta}}|[\beta^{*}(x_{0})]_{i}|$. When $\|z^{\prime}_{t+1}-z^{\prime}_{t}\|\leq\Delta^{\beta}/(6\sigma_{\beta}),\|A^{\prime}z^{\prime*}_{\delta}(x_{0},v_{t+1})-b\|\leq\Delta^{\beta}/(6\sigma_{z})$, we have
$\|v_{t+1}-v^{*}(x_{0})\|\leq$ $\sigma_{uzx0}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|+\sigma_{v2x0}\|A^{\prime}z^{\prime*}(x_{0},v_{t+1})-b\|+\sigma_{v1x0}\|A^{\prime}z^{\prime}_{t}-b\|$
where $\sigma_{uzx0}=\frac{1}{\eta_{z}}+\frac{1+\sigma_{\max}}{\eta_{z}s_{\min}^{\beta}}+\frac{l_{g,1}\sigma_{\beta}}{s_{\min}^{\beta}}$, $\sigma_{v2x0}=\frac{l_{g,1}\sigma_{z}}{s_{\min}^{\beta}}$, $\sigma_{v1x0}=\frac{\rho_{2}\sigma_{\max}(A)}{s_{\min}^{\beta}}$.
Proof.
Denote the active set of $y^{*}_{\delta}(x_{0})$ as $\mathcal{I}$, $\mathcal{J}=[d_{h}]/\mathcal{I}$. We have $\alpha^{*}_{\delta}(x_{0})]_{\mathcal{I}}=0$ and $\alpha^{*}_{\delta}(x_{0})]_{\mathcal{J}}<0$. Note that
$\|\alpha_{t}-\alpha^{*}_{\delta}(x_{0})\|\leq$ $\|y^{\prime}_{t}-y^{\prime*}_{\delta}(x_{0})\|\leq\|y_{t}-y^{*}_{\delta}(x_{0},u_{t+1})\|+\|y^{*}_{\delta}(x_{0})-y^{*}_{\delta}(x_{0},u_{t+1})\|$
$\leq$ $\sigma_{\alpha}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{y}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|$
Define $\Delta=\min_{i\in\mathcal{J}}|[\alpha^{*}_{\delta}(x_{0})]_{i}|$. When $\|y^{\prime}_{t+1}-y^{\prime}_{t}\|\leq\Delta/(6\sigma_{\alpha}),\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|\leq\Delta/(6\sigma_{y})$, we have
$\|\alpha_{t}-\alpha^{*}_{\delta}(x_{0})\|\leq\frac{1}{3}\Delta$
$\|\alpha_{t+1}-\alpha_{t}\|\leq\frac{1}{6}\Delta$
Thus, for $\mathcal{J}$, we have
$[\alpha_{t}]_{\mathcal{J}}<0$
$[\alpha_{t+1}]_{\mathcal{J}}<0$
Therefore, there are no projection in the update of $\alpha_{t+1}$ and we have
$\|[u_{t+1}]_{\mathcal{J}}-[u^{*}_{\delta}(x_{0})]_{\mathcal{J}}\|=\|[u_{t+1}]_{\mathcal{J}}\|=\|\frac{1}{\eta_{y}}([\alpha_{t+1}]_{\mathcal{J}}-[\alpha_{t}]_{\mathcal{J}})\|\leq\frac{1}{\eta_{y}}\|\alpha_{t+1}-\alpha_{t}\|$
Moreover, for $\mathcal{I}$, we have
$\nabla_{y}g_{\delta}(x_{0},y^{*}_{\delta}(x_{0}))+A_{\mathcal{I}}^{\top}[u^{*}_{\delta}(x_{0})]_{\mathcal{I}}=0$
Thus,
$\nabla_{y}g_{\delta}(x_{0},y_{t})+A_{\mathcal{I}}^{\top}[u_{t+1}]_{\mathcal{I}}+A_{\mathcal{J}}^{\top}[u_{t+1}]_{\mathcal{J}}+\rho_{1}A^{\top}(A^{\prime}y^{\prime}_{t}-b)=\frac{1}{\eta_{y}}(y_{t+1}-y_{t})$
Suppose $\sigma_{\min}(A_{\mathcal{I}})=s_{\min}$ , we have
$\|[u_{t+1}]_{\mathcal{I}}-[u^{*}_{\delta}(x_{0})]_{\mathcal{I}}\|$
$\leq$ $\frac{1}{s_{\min}}[\frac{1}{\eta_{y}}\|y_{t+1}-y_{t}\|+L_{g}\|y_{t}-y^{*}_{\delta}(x_{0})\|+\rho_{1}\sigma_{\max}(A)\|A^{\prime}y^{\prime}_{t}-b\|+\sigma_{\max}(A)\|u_{t+1}]_{\mathcal{J}}\|]$
$\leq$ $\frac{1}{s_{\min}}\Bigg[\left(\frac{1+\sigma_{\max}}{\eta_{y}}+L_{g}\sigma_{\alpha}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+L_{g}\sigma_{y}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|$
$+\rho_{1}\sigma_{\max}(A)\|A^{\prime}y^{\prime}_{t}-b\|\Bigg]$
Thus,
$\|u_{t+1}-u^{*}_{\delta}(x_{0})\|\leq$ $\sigma_{uyx0}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{u2x0}\|A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{t+1})-b\|+\sigma_{u1x0}\|A^{\prime}y^{\prime}_{t}-b\|$
where $\sigma_{uyx0}=\frac{1}{\eta_{y}}+\frac{1+\sigma_{\max}}{\eta_{y}s_{\min}}+\frac{L_{g}\sigma_{\alpha}}{s_{\min}}$, $\sigma_{u2x0}=\frac{L_{g}\sigma_{y}}{s_{\min}}$, $\sigma_{u1x0}=\frac{\rho_{1}\sigma_{\max}(A)}{s_{\min}}$
Similar conditions and conclusions also hold for $\|v_{t+1}-v^{*}(x_{0})\|$.
Appendix E Proofs of Theorem˜5.2 and Corollary˜5.3
In this section, we provide proofs for Theorem˜5.2 and Corollary˜5.3. We first introduce the additional notations and lemmas that will be used in this section.
Notations
$K(x,y^{\prime},z^{\prime},u,v)=\phi_{\delta}(x,y,z)+u^{\top}(Bx+A^{\prime}y^{\prime}-b)-v^{\top}(Bx+A^{\prime}z^{\prime}-b)$
$+\frac{\rho_{1}}{2}\|Bx+A^{\prime}y^{\prime}-b\|^{2}-\frac{\rho_{2}}{2}\|Bx+A^{\prime}z^{\prime}-b\|^{2}$
$L_{K}=l_{\delta}+2l_{g,1}+\max\{\rho_{1},\rho_{2}\}\max\{\sigma_{\max}^{2}(A^{\prime}),\sigma_{\max}^{2}(B),\sigma_{\max}(B)\sigma_{\max}(A^{\prime})\}$
$\mu_{y}=\min\{\mu_{g}-l_{\delta}-\rho_{1}\sigma_{\max}^{2}(A),\frac{\rho_{1}}{2}\}$
$\mu_{z}=\min\{\mu_{g}-\rho_{2}\sigma_{\max}^{2}(A),\frac{\rho_{2}}{2}\}$
$y^{*}_{\delta}(x)=\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)$
$z^{*}(x)=\mathop{\rm argmin}_{z\in\mathcal{Y}(x)}g(x,z)$
$u^{*}_{\delta}(x)=\mathop{\rm argmax}_{u\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)+u^{\top}(Bx+Ay-b)$
$v^{*}(x)=\mathop{\rm argmax}_{v\in\mathbb{R}_{+}}\min_{z\in\mathcal{Y}(x)}g(x,y)+v^{\top}(Bx+Az-b)$
$y^{\prime*}_{\delta}(x,u)=\mathop{\rm argmin}_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v)$
$z^{\prime*}(x,v)=\mathop{\rm argmax}_{z^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v)$
$[y^{*}_{\delta}(x,u)^{\top},\alpha^{*}_{\delta}(x,u)^{\top}]^{\top}=y^{\prime*}_{\delta}(x,u)$
$[z^{*}(x,v)^{\top},\beta^{*}(x,v)^{\top}]^{\top}=z^{\prime*}(x,v)$
$d(x,z^{\prime},u,v)=K(x,y^{\prime*}_{\delta}(x,u),z^{\prime},u,v)$
$q(x,v)=\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x,v))-v^{\top}(Bx+A^{\prime}z^{\prime*}(x,v)-b)-\frac{\rho_{2}}{2}\|Bx+A^{\prime}z^{\prime*}(x,v)-b\|^{2}$
$V_{t}=\frac{1}{4}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{t},v_{t})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
Lemma˜5.1
When the LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, the optimal Lagrangian multipliers of $y^{*}(x)$ and $y^{*}_{\delta}(x)$ are unique and we have
$u^{*}_{\delta}(x)$ $=\mathop{\rm argmax}_{u\in\mathbb{R}_{+}}\min_{y\in\mathcal{Y}(x)}g_{\delta}(x,y)+u^{\top}h(x,y)=\mathop{\rm argmax}_{u\in\mathbb{R}^{d_{h}}}\min_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v),$
$v^{*}(x)$ $=\mathop{\rm argmax}_{v\in\mathbb{R}_{+}}\min_{z\in\mathcal{Y}(x)}g(x,z)+v^{\top}h(x,z)=\mathop{\rm argmin}_{v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v).$
Proof.
Suppose
$v_{1}=\mathop{\rm argmax}_{v\in\mathbb{R}_{+}}\min_{z\in\mathcal{Y}(x)}g(x,y),$
$v_{2}=\mathop{\rm argmax}_{u\in\mathbb{R}^{d_{h}}}\min_{y^{\prime}\in\mathcal{P}_{y}}K(x,y^{\prime},z^{\prime},u,v).$
The KKT conditions for $v_{1}$ are
$\nabla_{y}g(x,z^{*}(x))+A^{\top}v_{1}=0$
$Bx+Az^{*}(x)-b\leq 0$
$v_{1}\geq 0$
$v_{1}^{\top}(Bx+Az^{*}(x)-b)=0$
The KKT conditions for $v_{2}$ are
$\nabla_{y}g(x,z^{*}(x))+A^{\top}v_{1}+\rho_{2}(Bx+A^{\prime}z^{\prime*}(x)-b)=0$
$Bx+A^{\prime}z^{\prime*}(x)-b=0$
$\beta\leq 0$
$v_{1}\in\partial I_{-}(\beta)$
Note that these two KKT conditions are equivalent. Moreover, since the LICQ condition (Definition˜4.3) holds for $z^{*}(x)$, we have $v_{1}=v_{2}$. Similar conditions and conclusions also hold for $u^{*}_{\delta}(x)$. ∎
Lemma E.1.
$K(x,y^{\prime},z^{\prime},u,v)$ is $\mu_{y}$-strongly w.r.t. $y^{\prime}$, $\mu_{z}$-strongly concave w.r.t. $z^{\prime}$, and $L_{K}$-smooth w.r.t. $x,y^{\prime},z^{\prime}$.
Proof.
According to Lemma˜A.4, we know that $K(x,y^{\prime},z^{\prime},u,v)$ is $\mu_{y}$-strongly convex w.r.t. $y^{\prime}$, $\mu_{z}$-strongly concave w.r.t. $z^{\prime}$. Moreover
$\nabla_{x}K(x,y^{\prime},z^{\prime},u,v)=$ $\nabla_{x}\phi_{\delta}(x,y,z)+B^{\top}(u-v)+\rho_{1}B^{\top}(Bx+A^{\prime}y^{\prime}-b)$
$-\rho_{2}B^{\top}(Bx+A^{\prime}z^{\prime}-b)$
$\nabla_{y}K(x,y^{\prime},z^{\prime},u,v)=$ $\nabla_{y}\phi_{\delta}(x,y,z)+A^{\top}u+\rho_{1}A^{\top}(A^{\prime}y^{\prime}-b)$
$\nabla_{z}K(x,y^{\prime},z^{\prime},u,v)=$ $\nabla_{z}\phi_{\delta}(x,y,z)-A^{\top}v-\rho_{2}A^{\top}(A^{\prime}z^{\prime}-b)$
Thus, $K(x,y^{\prime},z^{\prime},u,v)$ is $L_{K}$-smooth w.r.t. $x,y^{\prime},z^{\prime}$. ∎
Lemma E.2.
$\|u_{t+1}-u^{*}_{\delta}(x_{t})\|\leq\sigma_{uy}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{u1}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|+\sigma_{u2}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|$
$\|v_{t+1}-v^{*}(x_{t})\|\leq\sigma_{vz}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|+\sigma_{v1}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|+\sigma_{v2}\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|$
where $\sigma_{uy}=\frac{1}{\sigma_{\min}(A)}(\frac{1}{\eta_{y}}+\sigma_{\alpha}l_{g,1})$, $\sigma_{u1}=\rho_{1}$, $\sigma_{u2}=\frac{\sigma_{y}l_{g,1}}{\sigma_{\min}(A)}$, $\sigma_{vz}=\frac{1}{\sigma_{\min}(A)}(\frac{1}{\eta_{z}}+\sigma_{\beta}l_{g,1})$, $\sigma_{v1}=\rho_{2}$, $\sigma_{v2}=\frac{\sigma_{z}l_{g,1}}{\sigma_{\min}(A)}$
Proof.
By the optimality condition at $y^{*}_{\delta}(x_{t})$, we have
$\nabla_{y}g_{\delta}(x_{t},y^{*}_{\delta}(x_{t}))+A^{\top}u^{*}_{\delta}(x_{t})=0.$
The update rule of $y_{t}$:
$-\frac{y_{t+1}-y_{t}}{\eta_{y}}=\nabla_{y}g_{\delta}(x_{t},y_{t})+A^{\top}u_{t+1}+\rho_{1}A^{\top}(Bx_{t}+Ay_{t}-b-\alpha_{t}),$
Putting together, we have
$A^{\top}(u^{*}_{\delta}(x_{t})-u_{t+1})$
$=$ $\frac{y_{t+1}-y_{t}}{\eta_{y}}+\nabla_{y}g_{\delta}(x_{t},y_{t})-\nabla_{y}g_{\delta}(x_{t},y^{*}_{\delta}(x_{t}))+\rho_{1}A^{\top}(Bx_{t}+Ay_{t}-b-\alpha_{t})$
Since $A$ has full row rank, we have
$u^{*}_{\delta}(x_{t})-u_{t+1}=$ $(AA^{\top})^{-1}A\bigg[\frac{y_{t+1}-y_{t}}{\eta_{y}}+\nabla_{y}g_{\delta}(x_{t},y_{t})-\nabla_{y}g_{\delta}(x_{t},y^{*}_{\delta}(x_{t}))\bigg]$
$+\rho_{1}(Bx_{t}+Ay_{t}-b-\alpha_{t}),$
and
$\|u_{t+1}-u^{*}_{\delta}(x_{t})\|$
$\leq$ $\frac{1}{\sigma_{\min}(A)\eta_{y}}\|y_{t+1}-y_{t}\|+\frac{l_{g,1}}{\sigma_{\min}(A)}\|y_{t}-y^{*}_{\delta}(x_{t})\|+\rho_{1}\|Bx_{t}+Ay_{t}-b-\alpha_{t}\|$
Moreover,
$\|y_{t}-y^{*}_{\delta}(x_{t})\|\leq$ $\|y_{t}-y^{*}_{\delta}(x_{t},u_{t+1})\|+\|y^{*}_{\delta}(x_{t})-y^{*}_{\delta}(x_{t},u_{t+1})\|$
$\leq$ $\sigma_{\alpha}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{y}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|$
where the last equality is due to Lemma˜D.2.
Thus,
$\|u_{t+1}-u^{*}_{\delta}(x_{t})\|\leq\sigma_{uy}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|+\sigma_{u1}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|+\sigma_{u2}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|$
where $\sigma_{uy}=\frac{1}{\sigma_{\min}(A)}(\frac{1}{\eta_{y}}+\sigma_{\alpha}l_{g,1})$, $\sigma_{u1}=\rho_{1}$, $\sigma_{u2}=\frac{\sigma_{y}l_{g,1}}{\sigma_{\min}(A)}$.
Similarly, we have
$\|v_{t+1}-v^{*}(x_{t})\|\leq\sigma_{vz}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|+\sigma_{v1}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|+\sigma_{v2}\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|$
where $\sigma_{vz}=\frac{1}{\sigma_{\min}(A)}(\frac{1}{\eta_{z}}+\sigma_{\beta}l_{g,1})$, $\sigma_{v1}=\rho_{2}$, $\sigma_{v2}=\frac{\sigma_{z}l_{g,1}}{\sigma_{\min}(A)}$.
Lemma E.3.
$\|z^{\prime*}(x_{1})-z^{\prime*}(x_{2})\|\leq\sigma_{zb}\|x_{1}-x_{2}\|$
$\|y^{\prime*}_{\delta}(x_{1})-z^{\prime*}_{\delta}(x_{2})\|\leq\sigma_{yb}\|x_{1}-x_{2}\|$
$\|v^{*}(x_{1})-v^{*}(x_{2})\|\leq\sigma_{vb}\|x_{1}-x_{2}\|$
$\|u^{*}_{\delta}(x_{1})-u^{*}_{\delta}(x_{2})\|\leq\sigma_{ub}\|x_{1}-x_{2}\|$
where $\sigma_{zb}=\sigma_{\max}(A)\sigma_{z}+\sigma_{zx}$, $\sigma_{yb}=\sigma_{\max}(A)\sigma_{y}+\sigma_{yx}$, $\sigma_{vb}=\frac{l_{g,1}\sigma_{zb}}{\sigma_{\min}(A)}$, $\sigma_{ub}=\frac{L_{g}\sigma_{yb}}{\sigma_{\min}(A)}$.
Proof.
Here, we introduce an additional notation: $z^{\prime*}(x;w)=\mathop{\rm argmin}_{z^{\prime}\in\mathcal{G}(w)}g(x,z)+\frac{\rho_{2}}{2}\|Bx+A^{\prime}z^{\prime}-b\|^{2}$, where $\mathcal{G}(w)=\{z^{\prime}\in\mathcal{P}_{y}|Bw+A^{\prime}z^{\prime}-b=0\}$. We can notice that $z^{\prime*}(x;x)=z^{\prime*}(x)$. According to Lemma˜A.3, we have $\|z^{\prime*}(x_{1};x_{1})-z^{\prime*}(x_{1};x_{2})\|\leq\sigma_{z}\|Ax_{1}-Ax_{2}\|$. Moreover, we have $\|z^{\prime*}(x_{1};x_{2})-z^{\prime*}(x_{2};x_{2})\|\leq\sigma_{zx}\|x_{1}-x_{2}\|$. Thus, $\|z^{\prime*}(x_{1})-z^{\prime*}(x_{2})\|\leq[\sigma_{\max}(A)\sigma_{z}+\sigma_{zx}]\|x_{1}-x_{2}\|$. Moreover,
$\nabla_{y}g(x,z^{*}(x_{1}))+A^{\top}v^{*}(x_{1})=0$
$\nabla_{y}g(x,z^{*}(x_{2}))+A^{\top}v^{*}(x_{2})=0$
Thus,
$\|v^{*}(x_{1})-v^{*}(x_{2})\|\leq\frac{l_{g,1}\sigma_{zb}}{\sigma_{\min}(A)}\|x_{1}-x_{2}\|$
Similarly, we have $\|y^{\prime*}_{\delta}(x_{1})-y^{\prime*}_{\delta}(x_{2})\|\leq[\sigma_{\max}(A)\sigma_{y}+\sigma_{yx}]\|x_{1}-x_{2}\|$ and $\|u^{*}_{\delta}(x_{1})-u^{*}_{\delta}(x_{2})\|\leq\frac{L_{g}\sigma_{yb}}{\sigma_{\min}(A)}\|x_{1}-x_{2}\|$
E.1 Potential function
In this subsection, we will prove the following descent lemma for $V_{t}$.
Lemma E.4.
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$, $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y}=1/(4L_{K})$, $\eta_{z}=2/(L_{K}+4L_{d})$, $\eta_{x}=\min\{\eta_{y}\mu_{y}^{2}/(640L_{K}^{2}),\eta_{z}\mu_{z}^{2}/(640L_{K}^{2}),\eta_{u}/(240(\sigma_{y}^{2}+\sigma_{u2}^{2}+\sigma_{u1}^{2})L_{K}^{2}),\eta_{v}/(240(\sigma_{z}^{2}+\sigma_{v2}^{2}+\sigma_{v1}^{2})L_{K}^{2}),2/(L_{K}+4L_{d}+8L_{q}),1/(1920\eta_{y}L_{K}^{2}\sigma^{2}_{uy}),1/(1920\eta_{z}L_{K}^{2}\sigma^{2}_{uz})\}$, $\eta_{u}=\eta_{y}\mu_{y}^{2}/(256\sigma_{\max}^{2}(A))$, $\eta_{v}=\eta_{z}\mu_{z}^{2}/(256\sigma_{\max}^{2}(A))$, we have
$V_{t}-V_{t+1}\geq$ $\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{16\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\frac{1}{8\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|Bx+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}+\frac{\eta_{v}}{2}\|Bx+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}$
$+\frac{\eta_{u}}{4}\|Bx+A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{v}}{4}\|Bx+A^{\prime}z^{\prime}_{t}-b\|^{2}+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x_{t})\|^{2}$ (42)
Thus,
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}}(V_{0}-\min_{t}V_{t})$ (43)
Proof.
First, for function $d$, we have
$d(x_{t},z^{\prime}_{t},u_{t+1},v_{t+1})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
$=$ $K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t}),z^{\prime}_{t},u_{t},v_{t})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t},v_{t})$
$\geq$ $-\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{u}(Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b)^{\top}(Bx_{t}+A^{\prime}y^{\prime}_{t}-b)$
Note that
$\nabla_{x}d(x,z^{\prime},u,v)=$ $\nabla_{x}\phi_{\delta}(x,y^{*}_{\delta}(x,u),z)+B^{\top}(u-v)+\rho_{1}B^{\top}(Bx+A^{\prime}y^{\prime*}_{\delta}(x,u)-b)$
$-\rho_{2}B^{\top}(Bx+A^{\prime}z^{\prime}-b)$
$\nabla_{z^{\prime}}d(x,z^{\prime},u,v)=$ $\begin{bmatrix}\nabla_{z}g(x,z)-A^{\top}v-\rho_{2}A^{\top}(Bx+A^{\prime}z^{\prime}-b)\\ v+(Bx+A^{\prime}z^{\prime}-b)\end{bmatrix}$
Thus, according to Lemma˜D.2, we know that $\nabla_{x}d(x,z^{\prime},u,v)$ is $(L_{K}+L_{K}\sigma_{yx})$-continuous w.r.t. $x,z^{\prime}$ and $\nabla_{z}^{\prime}d(x,z^{\prime},u,v)$ is $L_{K}$-continuous w.r.t. $x,z^{\prime}$. Define $L_{d}=\max\{L_{K}+L_{K}\sigma_{yx},L_{K}\}$. We have
$d(x_{t+1},z^{\prime}_{t+1},u_{t+1},v_{t+1})-d(x_{t},z^{\prime}_{t},u_{t+1},v_{t+1})$
$\geq$ $\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),x_{t+1}-x_{t}\rangle+\langle\nabla_{z^{\prime}}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),z^{\prime}_{t+1}-z^{\prime}_{t}\rangle$
$-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
$\geq$ $\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),x_{t+1}-x_{t}\rangle+\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
Then, for function $q$, we have
$q(x_{t},v_{t})-q(x_{t},v_{t+1})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t}),u_{t},v_{t})-K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t+1})$
$\geq$ $K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t})-K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u_{t},v_{t+1})$
$\geq$ $\eta_{v}(Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b)^{\top}(Bx_{t}+A^{\prime}z^{\prime}_{t}-b)$
Note that
$\nabla_{x}q(x,v)=$ $\nabla_{x}\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x,v))+B^{\top}(u^{*}_{\delta}(x)-v)+\rho_{1}B^{\top}(Bx+A^{\prime}y^{\prime*}_{\delta}(x)-b)$
$-\rho_{2}B^{\top}(Bx+A^{\prime}z^{\prime*}(x,v)-b)$
Thus, according to Lemma˜D.2 and E.3, $q(\cdot,v)$ is $L_{q}=(L_{K}+L_{K}\sigma_{zx}+L_{K}\sigma_{yb}+\sigma_{\max}(B)\sigma_{ub})$-smooth. We have
$q(x_{t},v_{t+1})-q(x_{t+1},v_{t+1})$
$\geq$ $\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u^{*}(x_{t}),v_{t+1}),x_{t}-x_{t+1}\rangle-\frac{L_{q}}{2}(\|x_{t+1}-x_{t}\|^{2})$
Finally, for function $K$, we have
$K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})-K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})=-\eta_{u}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}+\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2},$
and
$K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})-K(x_{t+1},y^{\prime}_{t+1},z^{\prime}_{t+1},u_{t+1},v_{t+1})$
$\geq$ $\frac{1}{\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$-\frac{L_{K}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}).$
Thus, for $V_{t}$, we have
$V_{t}-V_{t+1}$
$\geq$ $\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1}),x_{t+1}-x_{t}\rangle+\frac{1}{\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\eta_{u}(Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b)^{\top}(Bx_{t}+A^{\prime}y^{\prime}_{t}-b)-\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}$
$-\frac{L_{d}}{2}(\|x_{t+1}-x_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})+2\langle\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u^{*}(x_{t}),v_{t+1}),x_{t}-x_{t+1}\rangle$
$+2\eta_{v}(Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b)^{\top}(Bx_{t}+A^{\prime}z^{\prime}_{t}-b)-L_{q}(\|x_{t+1}-x_{t}\|^{2})$
$+\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}+\frac{1}{4\eta_{y}}\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\frac{1}{4\eta_{z}}\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}-\frac{\eta_{u}}{4}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}$
$+\frac{\eta_{v}}{4}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}-\frac{L_{K}}{8}(\|x_{t+1}-x_{t}\|^{2}+\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}+\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2})$
$\geq$ $-\|\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t},u_{t+1}),z^{\prime}_{t},u_{t+1},v_{t+1})-\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u^{*}(x_{t}),v_{t+1})\|\|x_{t+1}-x_{t}\|$
$+\left(\frac{\eta_{u}}{2}-\frac{\eta_{u}}{4}\right)\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-\frac{\eta_{u}}{2}\|A^{\prime}y^{\prime}_{t}-A^{\prime}y_{\delta}^{\prime*}(x_{t},u_{t+1})\|^{2}$
$+\left(\eta_{v}+\frac{\eta_{v}}{4}-\eta_{v}\right)\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\eta_{v}\|A^{\prime}z^{\prime}_{t}-A^{\prime}z^{\prime*}(x_{t},v_{t+1})\|^{2}$
$-\frac{\eta_{x}}{2}\|\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u^{*}(x_{t}),v_{t+1})-\nabla_{x}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t+1},v_{t+1})\|^{2}$
$+\frac{1}{2\eta_{x}}\|x_{t}-x_{t+1}\|^{2}+\frac{\eta_{x}}{2}\|\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t},v_{t+1}),u^{*}(x_{t}),v_{t+1})\|^{2}$
$+\left(\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{\eta_{z}}-\frac{1}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2},$
and
$V_{t}-V_{t+1}$
$\geq$ $-\frac{1}{4\eta_{x}}\|x_{t+1}-x_{t}\|^{2}-3\eta_{x}L_{K}^{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime*}_{\delta}(x_{t})\|^{2}-3\eta_{x}L_{K}^{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$-3\eta_{x}L_{K}^{2}\|u_{t+1}-u^{*}(x_{t})\|^{2}$
$+\frac{\eta_{u}}{4}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-\frac{\eta_{u}\sigma_{\max}^{2}(A)}{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}$
$+\frac{\eta_{v}}{4}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\eta_{v}\sigma_{\max}^{2}(A)\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$-2\eta_{x}L_{K}^{2}\|y^{\prime*}_{\delta}(x_{t})-y^{\prime*}_{\delta}(x_{t},u_{t+1})\|^{2}-2\eta_{x}L_{K}^{2}\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}-2\eta_{x}L_{K}^{2}\|u_{t+1}-u^{*}(x_{t})\|^{2}$
$-2\eta_{x}L_{K}^{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}+\frac{1}{2\eta_{x}}\|x_{t}-x_{t+1}\|^{2}$
$+\frac{\eta_{x}}{4}\|\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t}),u^{*}(x_{t}),v^{*}(x_{t}))\|^{2}-\eta_{x}L_{K}^{2}\|z^{\prime*}(x_{t})-z^{\prime*}(x_{t},v_{t+1})\|^{2}$
$-\eta_{x}L_{K}^{2}\|v_{t+1}-v^{*}(x_{t})\|^{2}$
$+\left(\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2},$
and
$V_{t}-V_{t+1}$
$\geq$ $\left(\frac{1}{2\eta_{x}}+\frac{1}{4\eta_{x}}-\frac{1}{4\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}-\left(5\eta_{x}L_{K}^{2}+\frac{\eta_{u}\sigma_{\max}^{2}(A)}{2}\right)\|y^{\prime*}_{\delta}(x_{t},u_{t+1})-y^{\prime}_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}-\left(5\eta_{x}L_{K}^{2}+\eta_{v}\sigma_{\max}^{2}(A)\right)\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime}_{t}\|^{2}$
$+\frac{\eta_{u}}{4}\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}+\frac{\eta_{u}}{2}\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}-5\eta_{x}L_{K}^{2}\|y_{\delta}^{\prime*}(x_{t},u_{t+1})-y^{\prime*}_{\delta}(x_{t})\|^{2}$
$+\frac{\eta_{v}}{4}\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}+\eta_{v}\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}-\eta_{x}L_{K}^{2}\|z^{\prime*}(x_{t},v_{t+1})-z^{\prime*}(x_{t})\|^{2}$
$+\frac{\eta_{x}}{4}\|\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t}),u^{*}(x_{t}),v^{*}(x_{t}))\|^{2}-5\eta_{x}L_{K}^{2}\|u_{t+1}-u^{*}(x_{t})\|^{2}-\eta_{x}L_{K}^{2}\|v_{t+1}-v^{*}(x_{t})\|^{2}$
$\geq$ $\left(\frac{1}{2\eta_{x}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-L_{q}\right)\|x_{t+1}-x_{t}\|^{2}$
$+\left(\frac{1}{4\eta_{y}}-\frac{L_{K}}{8}-\frac{20\eta_{x}L_{K}^{2}+2\eta_{u}\sigma_{\max}^{2}(A)}{\mu_{y}^{2}\eta_{y}^{2}}-15\eta_{x}L_{K}^{2}\sigma^{2}_{uy}\right)\|y^{\prime}_{t+1}-y^{\prime}_{t}\|^{2}$
$+\left(\frac{3}{4\eta_{z}}-\frac{L_{K}}{8}-\frac{L_{d}}{2}-\frac{20\eta_{x}L_{K}^{2}+4\eta_{v}\sigma_{\max}^{2}(A)}{\mu_{z}^{2}\eta_{z}^{2}}-3\eta_{x}L_{K}^{2}\sigma^{2}_{vz}\right)\|z^{\prime}_{t+1}-z^{\prime}_{t}\|^{2}$
$+\left(\frac{\eta_{u}}{2}-5\sigma_{y}^{2}\eta_{x}L_{K}^{2}-15\eta_{x}L_{K}^{2}\sigma^{2}_{u2}\right)\|Bx_{t}+A^{\prime}y^{\prime*}_{\delta}(x_{t},u_{t+1})-b\|^{2}$
$+\left(\eta_{v}-\sigma_{z}^{2}\eta_{x}L_{K}^{2}-3\eta_{x}L_{K}^{2}\sigma^{2}_{v2}\right)\|Bx_{t}+A^{\prime}z^{\prime*}(x_{t},v_{t+1})-b\|^{2}$
$+\left(\frac{\eta_{u}}{4}-15\eta_{x}L_{K}^{2}\sigma^{2}_{u1}\right)\|Bx_{t}+A^{\prime}y^{\prime}_{t}-b\|^{2}+\left(\frac{\eta_{v}}{4}-3\eta_{x}L_{K}^{2}\sigma^{2}_{v1}\right)\|Bx_{t}+A^{\prime}z^{\prime}_{t}-b\|^{2}+\frac{\eta_{x}}{4}\|\nabla\phi_{\delta}(x)\|^{2}$
where the last equality is due to Lemma˜D.2, E.2 and $\nabla_{x}K(x_{t},y^{\prime*}_{\delta}(x_{t}),z^{\prime*}(x_{t}),u^{*}(x_{t}),v^{*}(x_{t}))=\nabla\phi_{\delta}(x)$.
Proof of Theorem˜5.2
When $\delta\leq\mu_{g}/(2l_{f,1})$, $0\leq\rho_{1}\leq\frac{\mu_{g}-\delta l_{f,1}}{\sigma_{\max}^{2}(A)}$, $0\leq\rho_{2}\leq\frac{\mu_{g}}{\sigma_{\max}^{2}(A)}$, $\eta_{y}=1/(4L_{K})$, $\eta_{z}=2/(L_{K}+4L_{d})$, $\eta_{x}=\min\{\eta_{y}\mu_{y}^{2}/(640L_{K}^{2}),\eta_{z}\mu_{z}^{2}/(640L_{K}^{2}),\eta_{u}/(240(\sigma_{y}^{2}+\sigma_{u2}^{2}+\sigma_{u1}^{2})L_{K}^{2}),\eta_{v}/(240(\sigma_{z}^{2}+\sigma_{v2}^{2}+\sigma_{v1}^{2})L_{K}^{2}),2/(L_{K}+4L_{d}+8L_{q}),1/(1920\eta_{y}L_{K}^{2}\sigma^{2}_{uy}),1/(1920\eta_{z}L_{K}^{2}\sigma^{2}_{uz})\}$, $\eta_{u}=\eta_{y}\mu_{y}^{2}/(256\sigma_{\max}^{2}(A))$, $\eta_{v}=\eta_{z}\mu_{z}^{2}/(256\sigma_{\max}^{2}(A))$, according to (43), we have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}}(V_{0}-\min_{t}V_{t})$ (44)
Note that
$V_{t}=$ $\frac{1}{4}K(x_{t},y^{\prime}_{t},z^{\prime}_{t},u_{t},v_{t})+2q(x_{t},v_{t})-d(x_{t},z^{\prime}_{t},u_{t},v_{t})$
$\geq$ $2q(x_{t},v_{t})-\frac{3}{4}d(x_{t},z^{\prime}_{t},u_{t},v_{t})\geq\frac{5}{4}q(x_{t},v_{t})\geq\frac{5}{4}\phi_{\delta}(x_{t})\geq\frac{5\delta\Phi^{*}}{4}-\frac{5\delta^{2}l^{2}_{f,0}}{8\mu_{g}}$
Therefore, when $\delta=\Theta(\epsilon)$, with $T=O(\epsilon^{-4})$, we have $t\in[T]$, such that $\|\nabla\Phi_{\delta}(x_{t})\|=\|\frac{1}{\delta}\nabla\phi_{\delta}(x_{t})\|\leq\epsilon$.
Moreover, if we have $x_{0},y_{0},z_{0},u_{0},v_{0}$ such that
$\|y_{0}-y^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|u_{0}-u^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|z_{0}-z^{*}(x_{0})\|\leq O(\delta)$
$\|v_{0}-v^{*}(x_{0})\|\leq O(\delta)$
Then, set $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$, we can easily prove that
$\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|\leq\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u^{*}_{\delta}(x_{0}))-b\|+\sigma_{\max}(A^{\prime})\sigma_{yu}\|u_{0}-u^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}y_{0}^{\prime}-b\|\leq\sigma_{\max}(A^{\prime})\|y^{\prime}_{0}-y^{\prime*}(x_{0})\|\leq O(\delta)$
and similarly,
$\|Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|\leq O(\delta).$
Then, we have
$V_{0}=$ $\frac{1}{4}[\delta f(x_{0},y_{0})+(g(x_{0},y_{0})-g(x_{0},z_{0}))+u_{0}^{\top}(Bx_{0}+A^{\prime}y_{0}^{\prime}-b)-v_{0}^{\top}(Bx_{0}+A^{\prime}z_{0}^{\prime}-b)$
$+\frac{\rho_{1}}{2}\|Bx_{0}+A^{\prime}y_{0}^{\prime}-b\|^{2}-\frac{\rho_{2}}{2}\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|^{2}]$
$+2[\delta f(x_{0},y^{*}_{\delta}(x_{0}))+(g(x_{0},y^{*}_{\delta}(x_{0}))-g(x_{0},z^{\prime*}(x_{0},v_{0})))-v_{0}^{\top}(Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b)$
$-\frac{\rho_{2}}{2}\|Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|^{2}]-[\delta f(x_{0},y^{\prime*}_{\delta}(x_{0},u_{0}))+(g(x_{0},y^{\prime*}_{\delta}(x_{0},u_{0}))-g(x_{0},z_{0}))$
$+u_{0}^{\top}(Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b)-v_{0}^{\top}(Bx_{0}+A^{\prime}z_{0}^{\prime}-b)$
$+\frac{\rho_{1}}{2}\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|^{2}-\frac{\rho_{2}}{2}\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|^{2}]$
$\leq$ $\frac{5\delta\Phi(x_{0})}{4}+O(1)l_{f,1}(\|y_{0}-y^{*}(x_{0})\|+\|y^{*}_{\delta}(x_{0})-y^{*}(x_{0})\|+\|y^{*}_{\delta}(x_{0},u_{0})-y^{*}(x_{0})\|)$
$+O(1)C_{g}(\|z_{0}-y_{0}\|+\|y^{*}_{\delta}(x_{0})-z^{*}(x_{0},v_{0})\|+\|y^{*}_{\delta}(x_{0},u_{0})-z_{0}\|)$
$+O(\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|+\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|^{2}+\|Bx_{0}+A^{\prime}y_{0}^{\prime}-b\|+\|Bx_{0}+A^{\prime}y_{0}^{\prime}-b\|^{2}$
$+\|Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|+\|Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|^{2}$
$+\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|+\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|^{2})$
$\leq$ $\frac{5\delta\Phi(x_{0})}{4}+O(\delta)$
where $G=\max\{g(x_{0},y_{0}),g(x_{0},z_{0}),g(x_{0},y^{*}_{\delta}(x_{0}),g(x_{0},z^{*}(x_{0},v_{0})),g(x_{0},y^{*}_{\delta}(x_{0},u_{0})\}$, $\mathcal{C}=\{y\in\mathbb{R}^{d_{y}}|g(x_{0},y)\leq G\}$, $C_{g}=\sup_{y\in\mathcal{C}}\nabla_{y}g(x_{0},y)$. Since $g(x_{0},y)$ is strongly convex w.r.t $y$, its sub-level set $\mathcal{C}$ is compact and convex. Moreover, since $g$ is Lipschitz smooth, its gradient in this compact set $\mathcal{C}$ is upper bounded by an $O(1)$ constant $C_{g}$.
We can notice that
$V_{0}-\min_{t}V_{t}\leq\frac{5\delta[\Phi(x_{0})-\Phi^{*}]}{4}+O(\delta)=O(\epsilon)$
and
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla\Phi_{\delta}(x_{t})\|^{2}=\frac{1}{T\delta^{2}}\sum_{t=0}^{T-1}\|\nabla\phi_{\delta}(x_{t})\|^{2}\leq\frac{4}{T\eta_{x}\delta^{2}}(V_{0}-\min_{t}V_{t})=O\left(\frac{\epsilon^{-1}}{T}\right).$
Therefore, we can find an $\epsilon$-stationary point of $\Phi_{\delta}(x)$ with a complexity of $O(\epsilon^{-3})$.
Proof of Corollary˜5.3
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [16], we can find $\hat{v}$, $\hat{z}$ such that
$\|v_{0}-v^{*}(x_{0})\|\leq\delta$
$\|z_{0}-z^{*}(x_{0})\|\leq\delta$
with a complexity of $O(\log(\epsilon^{-1}))$.
Set $y_{0}=z_{0}=\hat{z}$, $u_{0}=v_{0}=\hat{v}$, $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$. We have
$\|y_{0}-y^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}y^{\prime*}_{\delta}(x_{0},u_{0})-b\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}y_{0}^{\prime}-b\|\leq O(\delta)$
$\|u_{0}-u^{*}_{\delta}(x_{0})\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}z^{\prime*}(x_{0},v_{0})-b\|\leq O(\delta)$
$\|Bx_{0}+A^{\prime}z_{0}^{\prime}-b\|\leq O(\delta)$
with $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points and apply Algorithm˜1, according to Theorem˜5.2, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$.
Thus, the total complexity is $O(\epsilon^{-3}+\log(\epsilon^{-1}))=O(\epsilon^{-3})$.
Appendix F Detailed Experimental Settings and Additional Experiments
We adapt and modify the code from [16]. The experiments on the toy example and hyperparameter optimization for SVM are conducted on an AMD EPYC 9554 64-Core Processor. The experiments on transportation network design are conducted on an Intel(R) Xeon(R) Platinum 8375C CPU. For the toy example, we set the hyperparameters for our algorithm as $\delta=0.1,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.01,\rho_{1}=\rho_{2}=1$.
F.1 Hyperparameter optimization in SVM
Support Vector Machines (SVMs) construct a machine learning model by identifying the best possible hyperplane that maximizes the separation margin between data points from different classes. In a hard-margin SVM, no misclassification is allowed, ensuring that all samples are correctly classified. Conversely, soft-margin SVMs allow certain samples to be misclassified to accommodate cases where a perfect separation is not feasible. To achieve this, slack variables $\xi$ are introduced to quantify classification violations for each sample.
We consider the same problem formulation as in [16], which is
$\min_{c}\quad\mathcal{L}_{\mathcal{D}_{\text{val}}}(w^{*},b^{*})=\sum_{(z_{\text{val}},l_{\text{val}})\in\mathcal{D}_{\text{val}}}\exp\left(1-l_{\text{val}}\left(z_{\text{val}}^{\top}w^{*}+b^{*}\right)\right)+\frac{1}{2}\|c\|^{2}$
$w^{*},b^{*},\xi^{*}=\arg\min_{w,b,\xi}\quad\frac{1}{2}\|w\|^{2}$
$\text{s.t.}\quad l_{\text{tr},i}(z_{\text{tr},i}^{\top}w+b)\geq 1-\xi_{i},\quad$ $\forall i\in\{1,\dots,|\mathcal{D}_{\text{tr}}|\}$
$\xi_{i}\leq c_{i},\quad$ $\forall i\in\{1,\dots,|\mathcal{D}_{\text{tr}}|\}.$
where $\mathcal{D}_{\text{tr}}\triangleq\{(z_{\text{tr},i},l_{\text{tr},i})\}_{i=1}^{|\mathcal{D}_{\text{tr}}|}$ is the training dataset, $\mathcal{D}_{\text{val}}\triangleq\{(z_{\text{val},i},l_{\text{val},i})\}_{i=1}^{|\mathcal{D}_{\text{val}}|}$ is the validation dataset, with $z_{\text{tr},i}$ ($z_{\text{val},i}$) being the features of sample $i$ and $l_{\text{tr},i}$ ($l_{\text{val},i}$) being its corresponding labels. The hyperparameter $c_{i}$ is introduced to bound the soft margin violation $\xi_{i}$. Note that the LL problem is equivalent to
$w^{*},b^{*}=\arg\min_{w,b,\xi}\quad\frac{1}{2}\|w\|^{2}$
$\text{s.t.}\quad l_{\text{tr},i}(z_{\text{tr},i}^{\top}w+b)\geq 1-c_{i},\quad$ $\forall i\in\{1,\dots,|\mathcal{D}_{\text{tr}}|\}$
Then upper level variables are $c$, and the lower level variables are $w,b$. Thus, we have a coupled linear constraint for the lower level problem.
We compared our algorithm SFLCB with GAM [45], LV-HBA [49], BLOCC [16], and BiC-GAFFA [48] on the diabetes dataset. For GAM, we follow the same implementation approach and hyperparameters as [16], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [45]. For BLOCC and SFLCB, we set the LL objective as $\frac{1}{2}\|w\|^{2}+\frac{\mu}{2}b^{2}$, where $\mu=0.01$ serves as a regularization term to make the LL problem strongly convex w.r.t $w,b$. For LV-HBA, we set $\alpha=0.01,\gamma_{1}=0.1,\gamma_{2}=0.1,\eta=0.001$. For BLOCC, we set $\gamma=12,\eta=0.01,T=20,T_{y}=20,\eta_{1g}=0.001,\eta_{1F}=0.00001,\eta_{2g}=0.0001,\eta_{2F}=0.0001$. For BiC-GAFFA, we set $p=0.3,\gamma_{1}=10,\gamma_{2}=0.01,\eta=0.01,\alpha=0.001,\beta=0.001,c_{0}=10,R=10$. For our algorithm SFLCB, we set $\delta=0.01,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.001,\rho_{1}=\rho_{2}=0.01$. The hyperparameters of LV-HBA, BLOCC are the same as those used in [16] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [48]. The experiments are conducted across 10 different random train-validation-test splits, and the average results along with one standard deviation are reported in Figure˜2.
F.2 Transportation network design
We consider the same setting as in [16]. In this setting, we act as the operator to design a new network that connects a set of stations $S$. Passengers then decide whether to use this network based on their rational decisions (lower level). The objective is to maximize the operator’s benefit (upper level). The operator can select a set of potential links $\mathcal{A}\subseteq S\times S$, and for each link $(i,j)\in\mathcal{A}$, determine its capacity $x_{ij}$. A link is constructed if $x_{ij}>0$; a larger $x_{ij}$ attracts more travelers, generating more revenue but incurring higher construction costs $c_{ij}$. Passenger demand is defined over a set of origin-destination pairs $\mathcal{K}\subseteq S\times S$. For each $(o,d)\in\mathcal{K}$, there is a known traffic demand $w_{od}$ and existing travel times $t^{\rm ext}_{od}$. We assume a single existing network. The fraction of passengers choosing the new network for each $(o,d)$ pair is denoted by $y_{od}$, and $y^{ij}_{od}$ represents the proportion using link $(i,j)$.
We summarize the notation as follows. We keep the notation the same as in [16].
  • $x_{ij}\in\mathbb{R}_{+}$, the capacity of the new network for the link $(i,j)\in\mathcal{A}$.
  • $y^{od}\in[0,1]$, the proportion of passengers from $(o,d)\in\mathcal{K}$ choosing the new network for their travel.
  • $y_{ij}^{od}\in[0,1]$, the proportion of passengers from $(o,d)\in\mathcal{K}$ choosing the new network and use the link $(i,j)\in\mathcal{A}$
  • $x=\{x_{ij}\}_{\forall(i,j)\in\mathcal{A}}$ are the upper-level variables to be optimized.
  • $\mathcal{X}=\mathbb{R}^{|\mathcal{A}|}_{+}$ represents the domain of $x$.
  • $y=\{y^{od},\{y_{ij}^{od}\}_{\forall(i,j)\in\mathcal{A}}\}_{\forall(o,d)\in\mathcal{K}}$ are the lower-level variables to be optimized
  • $\mathcal{Y}=[\varepsilon,1-\varepsilon]^{|\mathcal{K}|}\times[\varepsilon,1-\varepsilon]^{|\mathcal{A}||\mathcal{K}|}$, where $\varepsilon$ is a small positive number, represents the domain of $y$.
  • $w^{od}$, the total estimated demand for $(o,d)\in\mathcal{K}$.
  • $m^{od}$, the revenue obtained by the operator from a passenger in $(o,d)\in\mathcal{K}$.
  • $c_{ij}$, the construction cost per passenger for link $(i,j)\in\mathcal{A}$.
  • $t_{ij}$, the travel time for link $(i,j)\in\mathcal{A}$.
  • $t_{\rm ext}^{od}$, travel time on the existing network for passengers in $(o,d)\in\mathcal{K}$.
  • $\omega_{t}<0$, the coefficient associated with the travel time for passengers.
With these notions, we can introduce the bilevel formulations of this problem. At the upper level, the network operator seeks to maximize the overall profit by attracting more passengers while minimizing construction costs; thus, its objective is
$\min_{x\in\mathcal{X}}f(x,y^{*}_{g}(x)):=-\Bigg(\underbrace{\sum_{\forall(o,d)\in\mathcal{K}}m^{od}y^{\rm od*}(x)}_{\text{profit}}-\underbrace{\sum_{\forall(i,j)\in\mathcal{A}}c_{ij}x_{ij}}_{cost}\Bigg),$ (45)
where $y^{\rm od*}(x)$ are optimal lower-level passenger flows. For the lower-level, the objective is defined as finding the flow variables that maximize passenger utility and minimize flow entropy cost.
$\min_{y\in\mathcal{Y}}g(x,y):=$ $-\Bigg(\underbrace{\sum_{(o,d)\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}w^{od}\omega_{t}t_{ij}y_{ij}^{od}+\sum_{(o,d)\in\mathcal{K}}w^{od}\omega_{t}t_{\rm ext}^{od}(1-y^{od})}_{\rm passengers~utility}$ (46)
$\qquad+\underbrace{\sum_{(o,d)\in\mathcal{K}}w^{od}y^{od}(\ln(y^{od})-1)+\sum_{(o,d)\in\mathcal{K}}w^{od}(1-y^{od})(\ln(1-y^{od})-1)}_{\rm flow~entropy~cost}\Bigg)$
s.t. $\;\;\sum_{\forall j|(i,j)\in\mathcal{A}}y_{ij}^{od}-\sum_{\forall j|(j,i)\in\mathcal{A}}y_{ji}^{od}=\begin{cases}y^{od}\;\quad\text{if}\;i=o\\ -y^{od}\;\ \text{if}\;i=d\\ 0\;\;\qquad\text{otherwise}\end{cases}\forall i,(o,d)\in\mathcal{S}\times\mathcal{K}$ (47)
$\qquad\qquad\;\;\;\;\;\;\sum_{\forall(o,d)\in\mathcal{K}}w^{od}y_{ij}^{od}\leq x_{ij}\qquad\qquad\;\;\forall(i,j)\in\mathcal{A}$ (48)
where (47) are the flow-conservation constraints and (48) are the capacity constraints.
We consider the same 3 nodes and 9 nodes settings as in [16] and use the same lower level feasibility criteria. The hyperparameter settings used in Figure˜3 are listed below. For the 3 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.1$, $\rho_{1}=\rho_{2}=1000$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-4$, and $T=30000$. For the 9 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.25$, $\rho_{1}=\rho_{2}=50$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-5$, and $T=300000$. For BLOCC, we used the same hyperparameters as those used in [16].

### v2

@@
 1 Introduction
-In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter, ; franceschi2018bilevel, ; mackay2019self, ), data hypercleaning (shaban2019truncated, ), meta learning (rajeswaran2019meta, ; ji2020convergence, ), reinforcement learning (sutton2018reinforcement, ; hong2023two, ) and neural architecture search (liu2018darts, ; liang2019darts+, ); see survey papers (zhang2024introduction, ; liu2021investigating, ; sinha2017review, ) for additional discussions.
-While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation, ; liu2020generic, ; ji2021bilevel, ; dagreou2022framework, ; hong2023two, ; kwon2023fully, ), studies focusing on constrained BLO problems are relatively limited.
+In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter; franceschi2018bilevel; mackay2019self), data hypercleaning (shaban2019truncated), meta learning (rajeswaran2019meta; ji2020convergence), reinforcement learning (sutton2018reinforcement; hong2023two) and neural architecture search (liu2018darts; liang2019darts+); see survey papers (zhang2024introduction; liu2021investigating; sinha2017review) for additional discussions.
+While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation; liu2020generic; ji2021bilevel; dagreou2022framework; hong2023two; kwon2023fully), studies focusing on constrained BLO problems are relatively limited.
 In this paper, we consider the following BLO problem where the lower-level (LL) problem has coupled constraints:
@@
 For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$.
-This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized, ), hyperparameter optimization of constrained learning problems (xu2023efficient, ) and adversarial training (zhang2022revisiting, ) and draw significant attentions (tsaknakis2022implicit, ; khanduri2023linearly, ; kornowski2024first, ).
-A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ).
-For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ).
-However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ).
+This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized), hyperparameter optimization of constrained learning problems (xu2023efficient) and adversarial training (zhang2022revisiting) and draw significant attentions (tsaknakis2022implicit; khanduri2023linearly; kornowski2024first).
+A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation; ji2021bilevel; ji2022will; chen2022single; khanduri2021near).
+For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating).
+However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating).
 The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
-Recently, some first-order methods (kwon2023penalty, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ; kornowski2024first, ) have been proposed for addressing constrained BLO problems.
+Recently, some first-order methods (kwon2023penalty; yao2024constrained; yao2024overcoming; jiang2024primal; kornowski2024first) have been proposed for addressing constrained BLO problems.
 Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem.
@@
 Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation.
-Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty, ; jiang2024primal, ) for coupled constraints $\mathcal{Y}(x)$.
+Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty; jiang2024primal) for coupled constraints $\mathcal{Y}(x)$.
 To address these limitations, in this paper, we establish a rigorous theoretical justification for this reformulation (3) and propose a single-loop Hessian-free algorithm for the linearly constrained cases.
@@
   • We establish a rigorous theoretical connection between the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ by proving the closeness of their values and derivatives under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$ with certain assumptions, which provides strong justifications for the reformulation (3).
-  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem.
-and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies).
+  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem, and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies).
 The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal.
   • Our experiments on hyperparameter optimization in the support vector machine (SVM) and transportation network design problems validate the practical effectiveness and efficiency of the proposed SFLCB algorithm.
-Table 1: Comparison of our paper with (kwon2023penalty, ; jiang2024primal, ; kornowski2024first, ).
+Table 1: Comparison of our paper with (kwon2023penalty; jiang2024primal).
 More detailed introductions and discussions of other related works can be found in Section˜2.
-Here, the “Complexity” for (kwon2023penalty, ; jiang2024primal, ) and SFLCB means the complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$; for (kornowski2024first, ), it means the complexity needed to achieve the $(\delta,\epsilon)$-Goldstein stationarity defined in kornowski2024first .
+Here, the “Complexity” means the iteration complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$ (3).
 Methods LL Constraint Complexity Loops
@@
 jiang2024primal $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
-kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(d\delta^{-1}\epsilon^{-3})$ double
-kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(\delta^{-1}\epsilon^{-4})$ double
 SFLCB (ours) $Bx+Ay-b\leq 0$, $A$ is full row rank $O(\epsilon^{-3})$ single
@@
 BLO without constraints.
-One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter, ).
-It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation, ).
-Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ; yang2023accelerating, ).
-Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient, ; grazzi2020iteration, ; liu2021towards, ; bolte2022automatic, ).
+One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter).
+It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation).
+Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel; ji2022will; chen2022single; khanduri2021near; yang2023accelerating).
+Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient; grazzi2020iteration; liu2021towards; bolte2022automatic).
 Recently, penalty-based methods have gained traction as a promising approach for solving BLO.
-Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value, ; mehra2021penalty, ; liu2022bome, ; kwon2023fully, ; shen2023penalty, ; gao2022value, ; lu2024slm, ).
+Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value; mehra2021penalty; liu2022bome; kwon2023fully; shen2023penalty; gao2022value; lu2024slm; lu2024first).
 BLO with constraints.
@@
 One is based on the implicit gradient method.
-Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly, ).
+Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly).
 tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm.
@@
 xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem.
-However, these implicit gradient-type algorithms (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
+However, these implicit gradient-type algorithms (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
 Another commonly used approach for solving constrained BLO problems is based on penalty reformulation.
-For example, yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function.
+For example, lu2024first reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions.
+yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function.
 However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly.
@@
 For $i\in[d_{h}]$, $h_{i}(x,y)$ is convex with respect to $y$, $h_{i},\nabla h_{i},\nabla^{2}h_{i}$ are respectively Lipschitz continuous with $l_{h,0},l_{h,1},l_{h,2}$ jointly over $\mathcal{X}\times\mathcal{Y}(x)$.
-Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; jiang2024barrier, ; kwon2023penalty, ) to ensure the existence of $\nabla\Phi(x)$.
+Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; jiang2024barrier; kwon2023penalty) to ensure the existence of $\nabla\Phi(x)$.
 Lemma 4.7 (Theorem 2 in xu2023efficient).
@@
 In this section, we focus on a special and important case where the LL constraints are h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b.
-This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized, ; khanduri2023linearly, ), adversarial training (zhang2022revisiting, ; khanduri2023linearly, ), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6).
-For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty, ; jiang2024primal, ).
+This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized; khanduri2023linearly), adversarial training (zhang2022revisiting; khanduri2023linearly), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6).
+For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty; jiang2024primal).
 First, we introduce the following slackness parameters $\alpha,\beta\in\mathbb{R}^{d_{h}}_{-}$ and define $y^{\prime}=(y^{\top},\alpha^{\top})^{\top}$, $z^{\prime}=(z^{\top},\beta^{\top})^{\top}$.
@@
 Corollary 5.3.
-When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA has full row rank, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply projected gradient descent (PGD) for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, we can find $\hat{v}$, $\hat{z}$ such that $\|\hat{v}-v^{*}(x_{0})\|\leq\delta$ and $\|\hat{z}-z^{*}(x_{0})\|\leq\delta$ with a complexity of $O(\log(\epsilon^{-1}))$.
+When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA has full row rank, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply projected gradient descent (PGD) for $\max_{v\in\mathbb{R}^{d_{h}}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},z)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, we can find $\hat{v}$, $\hat{z}$ such that $\|\hat{v}-v^{*}(x_{0})\|\leq\delta$ and $\|\hat{z}-z^{*}(x_{0})\|\leq\delta$ with a complexity of $O(\log(\epsilon^{-1}))$.
 Set $y_{0}=z_{0}=\hat{z}$, $u_{0}=v_{0}=\hat{v}$, $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$.
@@
 6.2 Hyperparameter optimization in SVM
-Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ).
+Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient; yao2024constrained; yao2024overcoming; jiang2024primal).
 Here we consider the same problem formulation as in jiang2024primal, which formulates this problem as a coupled linearly constrained BLO problem.
-We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient, ), LV-HBA (yao2024constrained, ), BLOCC (jiang2024primal, ), and BiC-GAFFA (yao2024overcoming, ) on the diabetes dataset (dua2017uci, ).
+We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient), LV-HBA (yao2024constrained), BLOCC (jiang2024primal), and BiC-GAFFA (yao2024overcoming) on the diabetes dataset (dua2017uci).
 Results are plotted in Figure˜2.
@@
 We consider the two synthetic networks of 3 and 9 nodes, same as those considered in jiang2024primal.
-We compare SFLCB with BLOCC (jiang2024primal, ).
+We compare SFLCB with BLOCC (jiang2024primal).
 Results are plotted in Figure˜3, which indicate that SFLCB significantly outperforms BLOCC on this network design task.
@@
 A promising direction for future research is to extend the current results to stochastic environments or more general constraint structures.
-Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal, ) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.
+Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.
 Acknowledgements
@@
 Appendix A Useful Lemmas
-Lemma A.1 (Lemma 12 in [3]).
+Lemma A.1 (Lemma 12 in [chen2021communication]).
 Suppose $f(\cdot)$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{X}$ is a convex closed set, and $\eta\leq 1/l$.
@@
 $\|x-x^{*}\|\leq\frac{2}{\mu\eta}\|x-x^{+}\|$
-Lemma A.2 (Lemma 23 in [24]).
+Lemma A.2 (Lemma 23 in [lin2020near]).
 Suppose for any fixed $y\in\mathcal{Y}$, $f(x,y)$ is $\mu$-strongly convex w.r.t.
@@
 $\|x^{*}(y_{1})-x^{*}(y_{2})\|\leq\frac{l}{\mu}\|y_{1}-y_{2}\|$
-Lemma A.3 (Theorem 4.1 in [50]).
+Lemma A.3 (Theorem 4.1 in [zhang2022global]).
 Suppose $f(x):\mathbb{R}^{d_{1}}\to\mathbb{R}$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{P}=\{x|Cx\leq e\},\mathcal{S}(r)=\{x\in\mathcal{P}|Ax-b=r\}$, where $C,A\in\mathbb{R}^{d_{2}\times d_{1}}$, $e,b,r\in\mathbb{R}^{d_{2}}$.
@@
 Proof.
-Similar results and proofs of Theorem˜4.1 can also be found in [21, 16]. For completeness, we also provide our proofs for Theorem˜4.1 here.
+Similar results and proofs of Theorem˜4.1 can also be found in [kwon2023penalty, jiang2024primal]. For completeness, we also provide our proofs for Theorem˜4.1 here.
 Note that $y_{\delta}^{*}(x)$ satisfy
@@
 Proof.
-The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [43].
-According to Lemmas 2 and 3 in [16].
+The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [wachsmuth2013licq].
+According to Lemmas 2 and 3 in [jiang2024primal].
 We know that when ˜3.1, 3.2, 3.3 hold, $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, defining $\psi(x)=g(x,y^{*}(x))$ and $\psi_{\delta}(x)=g_{\delta}(x,y^{*}_{\delta}(x))$, we have
@@
 Proof.
-According to Theorem 2 in [45], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
+According to Theorem 2 in [xu2023efficient], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
 Moreover, we can give the explicit expression of $\nabla\Phi(x)$.
@@
 Additionally, $\nabla_{y}\bar{h}(x,y^{*}(x))$ has full row rank.
-Thus, according to Lemma A.2 in [20], $H$ is invertible.
+Thus, according to Lemma A.2 in [kornowski2024first], $H$ is invertible.
 We can calculate $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ with the following equation.
 $\begin{bmatrix}\nabla y^{*}(x)\\ \nabla\bar{\lambda}^{*}(x)\end{bmatrix}=H^{-1}\begin{bmatrix}-\nabla^{2}_{xy}g(x,y^{*}(x))-\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ -\nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}$ (20)
-According to [9], we have
+According to [ghadimi2018approximation], we have
 $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x)).$
@@
 Proof of Corollary˜5.3
-When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [16], we can find $\hat{v}$, $\hat{z}$ such that
+When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [jiang2024primal], we can find $\hat{v}$, $\hat{z}$ such that
 $\|v_{0}-v^{*}(x_{0})\|\leq\delta$
@@
 Appendix F Detailed Experimental Settings and Additional Experiments
-We adapt and modify the code from [16].
+We adapt and modify the code from [jiang2024primal].
 The experiments on the toy example and hyperparameter optimization for SVM are conducted on an AMD EPYC 9554 64-Core Processor.
@@
 To achieve this, slack variables $\xi$ are introduced to quantify classification violations for each sample.
-We consider the same problem formulation as in [16], which is
+We consider the same problem formulation as in [jiang2024primal], which is
 $\min_{c}\quad\mathcal{L}_{\mathcal{D}_{\text{val}}}(w^{*},b^{*})=\sum_{(z_{\text{val}},l_{\text{val}})\in\mathcal{D}_{\text{val}}}\exp\left(1-l_{\text{val}}\left(z_{\text{val}}^{\top}w^{*}+b^{*}\right)\right)+\frac{1}{2}\|c\|^{2}$
@@
 Then upper level variables are $c$, and the lower level variables are $w,b$. Thus, we have a coupled linear constraint for the lower level problem.
-We compared our algorithm SFLCB with GAM [45], LV-HBA [49], BLOCC [16], and BiC-GAFFA [48] on the diabetes dataset.
-For GAM, we follow the same implementation approach and hyperparameters as [16], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [45].
+We compared our algorithm SFLCB with GAM [xu2023efficient], LV-HBA [yao2024constrained], BLOCC [jiang2024primal], and BiC-GAFFA [yao2024overcoming] on the diabetes dataset.
+For GAM, we follow the same implementation approach and hyperparameters as [jiang2024primal], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [xu2023efficient].
 For BLOCC and SFLCB, we set the LL objective as $\frac{1}{2}\|w\|^{2}+\frac{\mu}{2}b^{2}$, where $\mu=0.01$ serves as a regularization term to make the LL problem strongly convex w.r.t $w,b$.
@@
 For our algorithm SFLCB, we set $\delta=0.01,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.001,\rho_{1}=\rho_{2}=0.01$.
-The hyperparameters of LV-HBA, BLOCC are the same as those used in [16] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [48].
+The hyperparameters of LV-HBA, BLOCC are the same as those used in [jiang2024primal] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [yao2024overcoming].
 The experiments are conducted across 10 different random train-validation-test splits, and the average results along with one standard deviation are reported in Figure˜2.
 F.2 Transportation network design
-We consider the same setting as in [16].
+We consider the same setting as in [jiang2024primal].
 In this setting, we act as the operator to design a new network that connects a set of stations $S$.
@@
 The fraction of passengers choosing the new network for each $(o,d)$ pair is denoted by $y_{od}$, and $y^{ij}_{od}$ represents the proportion using link $(i,j)$.
-We summarize the notation as follows. We keep the notation the same as in [16].
+We summarize the notation as follows. We keep the notation the same as in [jiang2024primal].
   • $x_{ij}\in\mathbb{R}_{+}$, the capacity of the new network for the link $(i,j)\in\mathcal{A}$.
@@
 where (47) are the flow-conservation constraints and (48) are the capacity constraints.
-We consider the same 3 nodes and 9 nodes settings as in [16] and use the same lower level feasibility criteria.
+We consider the same 3 nodes and 9 nodes settings as in [jiang2024primal] and use the same lower level feasibility criteria.
 The hyperparameter settings used in Figure˜3 are listed below.
@@
 For the 9 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.25$, $\rho_{1}=\rho_{2}=50$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-5$, and $T=300000$.
-For BLOCC, we used the same hyperparameters as those used in [16].
+For BLOCC, we used the same hyperparameters as those used in [jiang2024primal].

### v3

@@
 1 Introduction
-In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter; franceschi2018bilevel; mackay2019self), data hypercleaning (shaban2019truncated), meta learning (rajeswaran2019meta; ji2020convergence), reinforcement learning (sutton2018reinforcement; hong2023two) and neural architecture search (liu2018darts; liang2019darts+); see survey papers (zhang2024introduction; liu2021investigating; sinha2017review) for additional discussions.
-While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation; liu2020generic; ji2021bilevel; dagreou2022framework; hong2023two; kwon2023fully), studies focusing on constrained BLO problems are relatively limited.
+In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (Pedregosa, 2016; Franceschi et al., 2018; MacKay et al., 2019), data hypercleaning (Shaban et al., 2019), meta learning (Rajeswaran et al., 2019; Ji et al., 2020), reinforcement learning (Sutton, 2018; Hong et al., 2023) and neural architecture search (Liu et al., 2018; Liang et al., 2019); see survey papers (Zhang et al., 2024; Liu et al., 2021a; Sinha et al., 2017) for additional discussions.
+While numerous works for unconstrained BLO problems have been proposed (Ghadimi and Wang, 2018; Liu et al., 2020; Ji et al., 2021; Dagréou et al., 2022; Hong et al., 2023; Kwon et al., 2023a), studies focusing on constrained BLO problems are relatively limited.
 In this paper, we consider the following BLO problem where the lower-level (LL) problem has coupled constraints:
@@
 For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$.
-This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized), hyperparameter optimization of constrained learning problems (xu2023efficient) and adversarial training (zhang2022revisiting) and draw significant attentions (tsaknakis2022implicit; khanduri2023linearly; kornowski2024first).
-A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation; ji2021bilevel; ji2022will; chen2022single; khanduri2021near).
-For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating).
-However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating).
+This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (Yang et al., 2022), hyperparameter optimization of constrained learning problems (Xu and Zhu, 2023) and adversarial training (Zhang et al., 2022) and draw significant attentions (Tsaknakis et al., 2022; Khanduri et al., 2023; Kornowski et al., 2024).
+A popular approach for solving unconstrained BLO is implicit gradient descent (Ghadimi and Wang, 2018; Ji et al., 2021, 2022; Chen et al., 2022; Khanduri et al., 2021).
+For constrained BLO, several studies have extended this approach to accommodate different constraint settings (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023).
+However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023).
 The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
-Recently, some first-order methods (kwon2023penalty; yao2024constrained; yao2024overcoming; jiang2024primal; kornowski2024first) have been proposed for addressing constrained BLO problems.
+Recently, some first-order methods (Kwon et al., 2023b; Yao et al., 2024b, a; Jiang et al., 2024a; Kornowski et al., 2024) have been proposed for addressing constrained BLO problems.
 Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem.
-For example, yao2024constrained; yao2024overcoming reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem.
-However, neither yao2024constrained nor yao2024overcoming establishes clear relationships between the stationary points of their approximated problems and the original one.
-Works most closely related to ours are those by kwon2023penalty and jiang2024primal, both of which reformulated the problem (1) as
+For example, Yao et al.
+(2024b, a) reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem.
+However, neither Yao et al.
+(2024b) nor Yao et al.
+(2024a) establishes clear relationships between the stationary points of their approximated problems and the original one.
+Works most closely related to ours are those by Kwon et al. (2023b) and Jiang et al. (2024a), both of which reformulated the problem (1) as
 $\min_{x\in\mathcal{X},y\in\mathcal{Y}(x)}f(x,y)\quad\text{s.t. }g(x,y)-\min_{z\in\mathcal{Y}(x)}g(x,z)\leq 0,$ (2)
@@
 where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$.
-Based on this reformulation, kwon2023penalty proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and jiang2024primal proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
-However, kwon2023penalty only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration.
-Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation.
-Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty; jiang2024primal) for coupled constraints $\mathcal{Y}(x)$.
+Based on this reformulation, Kwon et al.
+(2023b) proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and Jiang et al.
+(2024a) proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
+However, Kwon et al.
+(2023b) only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration.
+Algorithms in Jiang et al.
+(2024a) require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation.
+Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (Kwon et al., 2023b; Jiang et al., 2024a) for coupled constraints $\mathcal{Y}(x)$.
 To address these limitations, in this paper, we establish a rigorous theoretical justification for this reformulation (3) and propose a single-loop Hessian-free algorithm for the linearly constrained cases.
@@
   • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem, and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies).
-The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal.
+The simple single-loop structure also makes our algorithm easier to implement in practice compared to Jiang et al.
+(2024a).
   • Our experiments on hyperparameter optimization in the support vector machine (SVM) and transportation network design problems validate the practical effectiveness and efficiency of the proposed SFLCB algorithm.
-Table 1: Comparison of our paper with (kwon2023penalty; jiang2024primal).
+Table 1: Comparison of our paper with (Kwon et al., 2023b; Jiang et al., 2024a).
 More detailed introductions and discussions of other related works can be found in Section˜2.
@@
 Methods LL Constraint Complexity Loops
-kwon2023penalty $y\in\mathcal{Y}$, $\mathcal{Y}$ is a convex and compact set $O(\epsilon^{-3}\log(\epsilon^{-1}))$ single/double
-jiang2024primal $h(x,y)\leq 0$, LICQ holds $O(\epsilon^{-5}\log(\epsilon^{-1}))$ triple
-jiang2024primal $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
+Kwon et al. (2023b) $y\in\mathcal{Y}$, $\mathcal{Y}$ is a convex and compact set $O(\epsilon^{-3}\log(\epsilon^{-1}))$ single/double
+Jiang et al. (2024a) $h(x,y)\leq 0$, LICQ holds $O(\epsilon^{-5}\log(\epsilon^{-1}))$ triple
+Jiang et al. (2024a) $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
 SFLCB (ours) $Bx+Ay-b\leq 0$, $A$ is full row rank $O(\epsilon^{-3})$ single
@@
 BLO without constraints.
-One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter).
-It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation).
-Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel; ji2022will; chen2022single; khanduri2021near; yang2023accelerating).
-Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient; grazzi2020iteration; liu2021towards; bolte2022automatic).
+One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (Pedregosa, 2016).
+It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (Ghadimi and Wang, 2018).
+Later works improved the convergence rates and studied the gradient descent methods under various settings (Ji et al., 2021, 2022; Chen et al., 2022; Khanduri et al., 2021; Yang et al., 2023).
+Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (Maclaurin et al., 2015; Grazzi et al., 2020; Liu et al., 2021c; Bolte et al., 2022).
 Recently, penalty-based methods have gained traction as a promising approach for solving BLO.
-Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value; mehra2021penalty; liu2022bome; kwon2023fully; shen2023penalty; gao2022value; lu2024slm; lu2024first).
+Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (Liu et al., 2021b; Mehra and Hamm, 2021; Liu et al., 2022; Kwon et al., 2023a; Shen and Chen, 2023; Gao et al., 2022; Lu, 2024; Lu and Mei, 2024).
 BLO with constraints.
@@
 One is based on the implicit gradient method.
-Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly).
-tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm.
-khanduri2023linearly proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$.
-xu2023efficient used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$.
+Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (Khanduri et al., 2023).
+Tsaknakis et al.
+(2022) proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm.
+Khanduri et al.
+(2023) proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$.
+Xu and Zhu (2023) used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$.
 However, they only provided an asymptotic convergence analysis of their algorithm.
-xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem.
-However, these implicit gradient-type algorithms (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
+Xiao et al.
+(2023) proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem.
+However, these implicit gradient-type algorithms (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
 Another commonly used approach for solving constrained BLO problems is based on penalty reformulation.
-For example, lu2024first reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions.
-yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function.
+For example, Lu and Mei (2024) reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions.
+Yao et al.
+(2024b) reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function.
 However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly.
-yao2024overcoming reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm.
-Compared to yao2024constrained, yao2024overcoming did not need the projection operator to the coupled constraint set.
-However, both yao2024constrained and yao2024overcoming did not establish very clear relationships between the stationary points of their approximated problems and the original one.
-For example, yao2024overcoming only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one.
-Recently, tsaknakis2023implicit; jiang2024barrier proposed algorithms based on barrier approximation approach for constrained BLO problems.
+Yao et al.
+(2024a) reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm.
+Compared to Yao et al.
+(2024b), Yao et al.
+(2024a) did not need the projection operator to the coupled constraint set.
+However, both Yao et al.
+(2024b) and Yao et al.
+(2024a) did not establish very clear relationships between the stationary points of their approximated problems and the original one.
+For example, Yao et al.
+(2024a) only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one.
+Recently, Tsaknakis et al.
+(2023); Jiang et al.
+(2024b) proposed algorithms based on barrier approximation approach for constrained BLO problems.
 However, their algorithms also require the computations for the Hessian matrix.
-kwon2023penalty and jiang2024primal considered the same reformulation as ours.
-kwon2023penalty studied the case where the LL variables $y\in\mathcal{Y}$ are independent of $x$ and characterized the conditions under which the values and derivatives of $\Phi$ and $\Phi_{\delta}$ can be $O(\delta)$-close for $y\in\mathcal{Y}$ constraints.
-Compared with kwon2023penalty, we prove similar results under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
-Moreover, the algorithms in kwon2023penalty require the implementation of the projection operator to $\mathcal{Y}$ at each iteration, which can be costly.
-jiang2024primal studied the coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
-While jiang2024primal considered more general constraints than ours, however, it did not characterize the gap between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$.
+Kwon et al.
+(2023b) and Jiang et al.
+(2024a) considered the same reformulation as ours.
+Kwon et al.
+(2023b) studied the case where the LL variables $y\in\mathcal{Y}$ are independent of $x$ and characterized the conditions under which the values and derivatives of $\Phi$ and $\Phi_{\delta}$ can be $O(\delta)$-close for $y\in\mathcal{Y}$ constraints.
+Compared with Kwon et al.
+(2023b), we prove similar results under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
+Moreover, the algorithms in Kwon et al.
+(2023b) require the implementation of the projection operator to $\mathcal{Y}$ at each iteration, which can be costly.
+Jiang et al.
+(2024a) studied the coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
+While Jiang et al.
+(2024a) considered more general constraints than ours, however, it did not characterize the gap between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$.
 Our Theorem˜4.9 provides further justifications for their reformulation in coupled constraints.
-Moreover, compared with the double- and triple-loop algorithms in kwon2023penalty, we propose a single-loop algorithm SFLCB and prove an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$.
-Recently, kornowski2024first also proposed first-order methods for linearly constrained BLO.
+Moreover, compared with the double- and triple-loop algorithms in Kwon et al.
+(2023b), we propose a single-loop algorithm SFLCB and prove an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$.
+Recently, Kornowski et al.
+(2024) also proposed first-order methods for linearly constrained BLO.
 Especially, they proved a nearly optimal convergence rate $\tilde{O}(\epsilon^{-2})$ for linear equality constraints and proposed algorithms that can attain $(\delta,\epsilon)$-Goldstein stationarity for linear inequality constraints.
 However, their convergence rates for linear inequality constraints either have additional dependence on dimension $d$ (such as $\tilde{O}(d\delta^{-1}\epsilon^{-3})$) or need additional assumptions to access the exact optimal dual variable (such as $\tilde{O}(\delta^{-1}\epsilon^{-4})$), while we do not require the exact optimal dual variable assumption.
-Compared with the double-loop algorithms in kornowski2024first, our proposed single-loop one is easier to implement in practice.
+Compared with the double-loop algorithms in Kornowski et al.
+(2024), our proposed single-loop one is easier to implement in practice.
 Moreover, our techniques are also different from theirs under linear inequality constraints, thereby highlighting the distinct contributions and independent interests of our work.
@@
 $\phi_{\delta}(x)=\phi_{\delta}(x,y^{*}_{\delta}(x),z^{*}(x)).$
-Similar to Theorem 3.8 in kwon2023penalty, we have the following theorem to bound the difference between $\Phi$ and $\Phi_{\delta}$, as well as $y^{*}(x)$ and $y_{\delta}^{*}(x)$ in the coupled constraints.
+Similar to Theorem 3.8 in Kwon et al.
+(2023b), we have the following theorem to bound the difference between $\Phi$ and $\Phi_{\delta}$, as well as $y^{*}(x)$ and $y_{\delta}^{*}(x)$ in the coupled constraints.
 Theorem 4.1.
@@
 We say a point $y$ satisfy the LICQ condition if, for all $i\in\mathcal{I}_{y}$, $\nabla_{y}h_{i}(x,y)$ are linearly independent.
-Then, similar to Lemmas 2 and 3 in jiang2024primal, we have the following lemma.
+Then, similar to Lemmas 2 and 3 in Jiang et al. (2024a), we have the following lemma.
 Lemma 4.4.
@@
 While the gradients of $\Phi_{\delta}(x)$ exists under LICQ conditions, for general problem (1), $\Phi(x)$ is not guaranteed to be differentiable.
-For example, khanduri2023linearly provides an example where the LICQ condition holds, $\Phi(x)$ is non-differentiable at some points.
+For example, Khanduri et al.
+(2023) provides an example where the LICQ condition holds, $\Phi(x)$ is non-differentiable at some points.
 However, if a given $x$ satisfies the following conditions, then $\nabla\Phi(x)$ exists at $x$.
@@
 For $i\in[d_{h}]$, $h_{i}(x,y)$ is convex with respect to $y$, $h_{i},\nabla h_{i},\nabla^{2}h_{i}$ are respectively Lipschitz continuous with $l_{h,0},l_{h,1},l_{h,2}$ jointly over $\mathcal{X}\times\mathcal{Y}(x)$.
-Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; jiang2024barrier; kwon2023penalty) to ensure the existence of $\nabla\Phi(x)$.
-Lemma 4.7 (Theorem 2 in xu2023efficient).
+Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Jiang et al., 2024b; Kwon et al., 2023b) to ensure the existence of $\nabla\Phi(x)$.
+Lemma 4.7 (Theorem 2 in Xu and Zhu (2023)).
 When ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and LICQ (Definition˜4.3) condition hold for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
@@
 Thus, we expect that for $t\in[0,\delta]$, $y^{*}_{t}(x)$ will have similar properties as $y^{*}(x)$.
-Similar assumptions have also been used in kwon2023penalty to establish the non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
+Similar assumptions have also been used in Kwon et al.
+(2023b) to establish the non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
 Theorem 4.9.
@@
 $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|\leq O(\delta).$
-Similar non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$ has been established in kwon2023penalty; however, their bound is established only for the LL constraints $\mathcal{Y}$ that do not depend on $x$.
+Similar non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$ has been established in Kwon et al.
+(2023b); however, their bound is established only for the LL constraints $\mathcal{Y}$ that do not depend on $x$.
 Our Theorem˜4.9 provides a more general theoretical justification for the validity of the reformulation (3) for coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.
@@
 In this section, we focus on a special and important case where the LL constraints are h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b.
-This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized; khanduri2023linearly), adversarial training (zhang2022revisiting; khanduri2023linearly), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6).
-For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty; jiang2024primal).
+This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (Yang et al., 2022; Khanduri et al., 2023), adversarial training (Zhang et al., 2022; Khanduri et al., 2023), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6).
+For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (Kwon et al., 2023b; Jiang et al., 2024a).
 First, we introduce the following slackness parameters $\alpha,\beta\in\mathbb{R}^{d_{h}}_{-}$ and define $y^{\prime}=(y^{\top},\alpha^{\top})^{\top}$, $z^{\prime}=(z^{\top},\beta^{\top})^{\top}$.
@@
 $L_{\delta}(x,y^{\prime},z^{\prime},u,v)=$ $\phi_{\delta}(x,y,z)+u^{\top}(h(x,y)-\alpha)-v^{\top}(h(x,z)-\beta).$
-According to Proposition 5.3.4 in bertsekas2009convex, we know that
+According to Proposition 5.3.4 in Bertsekas (2009), we know that
 $\phi_{\delta}(x)=\min_{y^{\prime}\in\mathcal{P}_{y},v\in\mathbb{R}^{d_{h}}}\max_{z^{\prime}\in\mathcal{P}_{y},u\in\mathbb{R}^{d_{h}}}L_{\delta}(x,y^{\prime},z^{\prime},u,v).$
@@
 Then, we present the convergence results of SFLCB (Algorithm˜1) for coupled constraints h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $A$ is full row rank.
-Note that BLOCC in jiang2024primal that achieves the complexity of $O(\epsilon^{-3}\log(\epsilon^{-1}))$ also needs the matrix $A$ to be full row rank (See Table˜1).
+Note that BLOCC in Jiang et al.
+(2024a) that achieves the complexity of $O(\epsilon^{-3}\log(\epsilon^{-1}))$ also needs the matrix $A$ to be full row rank (See Table˜1).
 For $A$ that is not full row rank and $B=0$, we provide the convergence results in Theorem˜5.4 and Corollary˜5.5.
@@
 The full row rank property of $A$ is used in Lemma˜E.2 to bound $\|u_{t+1}-u^{*}_{\delta}(x_{t})\|$ and $\|v_{t+1}-v^{*}(x_{t})\|$.
-Since $A$ has full row rank, then according to Theorem 6 in jiang2024primal, we can easily find initial points satisfying (8)-(9) with a complexity of $O(\log(\epsilon^{-1}))$ and we have the following corollary.
+Since $A$ has full row rank, then according to Theorem 6 in Jiang et al.
+(2024a), we can easily find initial points satisfying (8)-(9) with a complexity of $O(\log(\epsilon^{-1}))$ and we have the following corollary.
 Corollary 5.3.
@@
 The proof of Corollary˜5.3 can be found in Appendix˜E.
-Thus, compared to jiang2024primal, we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ for the coupled linear constraint (See Table˜1).
+Thus, compared to Jiang et al.
+(2024a), we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ for the coupled linear constraint (See Table˜1).
 Additionally, we have the following convergence results for constraints h(y)=Aybh(y)=Ay-b, where $A$ is not required to have a full row rank.
@@
 Note that in Corollary˜5.5, we only need the LICQ condition holds for the initial $y^{*}(x_{0})$ and $y^{*}_{\delta}(x_{0})$, and we can achieve a total complexity of $O(\epsilon^{-3})$.
-Compared to the decoupled constrained setting in kwon2023penalty, we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ (See Table˜1).
+Compared to the decoupled constrained setting in Kwon et al.
+(2023b), we achieve an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ (See Table˜1).
 6 Experiments
@@
 Figure 1: Toy example.
-We consider the same constrained BLO problem that was studied in jiang2024primal, which is:
+We consider the same constrained BLO problem that was studied in Jiang et al. (2024a), which is:
 $\min_{x\in[0,3]}f(x,y^{*}(x))=\frac{e^{-y_{g}^{*}(x)+2}}{2+\cos(6x)}+\frac{1}{2}\ln\left((4x-2)^{2}+1\right)$ (12)
@@
 6.2 Hyperparameter optimization in SVM
-Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient; yao2024constrained; yao2024overcoming; jiang2024primal).
-Here we consider the same problem formulation as in jiang2024primal, which formulates this problem as a coupled linearly constrained BLO problem.
-We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient), LV-HBA (yao2024constrained), BLOCC (jiang2024primal), and BiC-GAFFA (yao2024overcoming) on the diabetes dataset (dua2017uci).
+Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (Xu and Zhu, 2023; Yao et al., 2024b, a; Jiang et al., 2024a).
+Here we consider the same problem formulation as in Jiang et al.
+(2024a), which formulates this problem as a coupled linearly constrained BLO problem.
+We conduct experiments comparing our SFLCB algorithm with GAM (Xu and Zhu, 2023), LV-HBA (Yao et al., 2024b), BLOCC (Jiang et al., 2024a), and BiC-GAFFA (Yao et al., 2024a) on the diabetes dataset (Dua et al., 2017).
 Results are plotted in Figure˜2.
@@
 6.3 Transportation network design
-We further conduct experiments on a transportation network design problem, following the same setting as in jiang2024primal.
+We further conduct experiments on a transportation network design problem, following the same setting as in Jiang et al.
+(2024a).
 In this setting, we act as the operator, whose profit serves as the upper-level objective and is influenced by passenger behavior, which is modeled in the lower-level problem.
 Detailed formulations and settings can be found in Appendix˜F.
-We consider the two synthetic networks of 3 and 9 nodes, same as those considered in jiang2024primal.
-We compare SFLCB with BLOCC (jiang2024primal).
+We consider the two synthetic networks of 3 and 9 nodes, same as those considered in Jiang et al.
+(2024a).
+We compare SFLCB with BLOCC (Jiang et al., 2024a).
 Results are plotted in Figure˜3, which indicate that SFLCB significantly outperforms BLOCC on this network design task.
@@
 In this paper, for coupled constrained BLO problem in Equation˜1, we theoretically analyzed the relationship between the original hyper-objective $\Phi$ and the reformulated function $\Phi_{\delta}$ in Equation˜3, providing a solid justification for the validity of the reformulation.
-Especially, for the linearly constrained case, we proposed SFLCB, a single-loop, Hessian-free algorithm, improving the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ over previous works kwon2023penalty; jiang2024primal.
+Especially, for the linearly constrained case, we proposed SFLCB, a single-loop, Hessian-free algorithm, improving the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ over previous works Kwon et al.
+(2023b); Jiang et al.
+(2024a).
 Our experiments on hyperparameter optimization for SVM and the transportation network design problem validated the practical efficiency of the proposed SFLCB algorithm.
@@
 A promising direction for future research is to extend the current results to stochastic environments or more general constraint structures.
-Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.
+Moreover, since the best-known complexity for first-order methods in unconstrained BLO (Huang, 2024) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.
 Acknowledgements
 The work of Wei Shen and Cong Shen was supported in part by the US National Science Foundation (NSF) under awards 2143559 and 2332060.
-The work of Jiawei Zhang was supported by the startup funding from the University of Wisconsin-Madison.
+The work of Jiawei Zhang was supported by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation.
 Appendix
@@
 Appendix A Useful Lemmas
-Lemma A.1 (Lemma 12 in [chen2021communication]).
+Lemma A.1 (Lemma 12 in [3]).
 Suppose $f(\cdot)$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{X}$ is a convex closed set, and $\eta\leq 1/l$.
@@
 $\|x-x^{*}\|\leq\frac{2}{\mu\eta}\|x-x^{+}\|$
-Lemma A.2 (Lemma 23 in [lin2020near]).
+Lemma A.2 (Lemma 23 in [24]).
 Suppose for any fixed $y\in\mathcal{Y}$, $f(x,y)$ is $\mu$-strongly convex w.r.t.
@@
 $\|x^{*}(y_{1})-x^{*}(y_{2})\|\leq\frac{l}{\mu}\|y_{1}-y_{2}\|$
-Lemma A.3 (Theorem 4.1 in [zhang2022global]).
+Lemma A.3 (Theorem 4.1 in [51]).
 Suppose $f(x):\mathbb{R}^{d_{1}}\to\mathbb{R}$ is $l$-smooth and $\mu$-strongly convex, $\mathcal{P}=\{x|Cx\leq e\},\mathcal{S}(r)=\{x\in\mathcal{P}|Ax-b=r\}$, where $C,A\in\mathbb{R}^{d_{2}\times d_{1}}$, $e,b,r\in\mathbb{R}^{d_{2}}$.
@@
 Proof.
-Similar results and proofs of Theorem˜4.1 can also be found in [kwon2023penalty, jiang2024primal]. For completeness, we also provide our proofs for Theorem˜4.1 here.
+Similar results and proofs of Theorem˜4.1 can also be found in [22, 16]. For completeness, we also provide our proofs for Theorem˜4.1 here.
 Note that $y_{\delta}^{*}(x)$ satisfy
@@
 Proof.
-The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [wachsmuth2013licq].
-According to Lemmas 2 and 3 in [jiang2024primal].
+The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [44].
+According to Lemmas 2 and 3 in [16].
 We know that when ˜3.1, 3.2, 3.3 hold, $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, defining $\psi(x)=g(x,y^{*}(x))$ and $\psi_{\delta}(x)=g_{\delta}(x,y^{*}_{\delta}(x))$, we have
@@
 Proof.
-According to Theorem 2 in [xu2023efficient], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
+According to Theorem 2 in [46], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
 Moreover, we can give the explicit expression of $\nabla\Phi(x)$.
@@
 Additionally, $\nabla_{y}\bar{h}(x,y^{*}(x))$ has full row rank.
-Thus, according to Lemma A.2 in [kornowski2024first], $H$ is invertible.
+Thus, according to Lemma A.2 in [20], $H$ is invertible.
 We can calculate $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ with the following equation.
 $\begin{bmatrix}\nabla y^{*}(x)\\ \nabla\bar{\lambda}^{*}(x)\end{bmatrix}=H^{-1}\begin{bmatrix}-\nabla^{2}_{xy}g(x,y^{*}(x))-\nabla^{2}_{xy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\\ -\nabla_{x}\bar{h}(x,y^{*}(x))\end{bmatrix}$ (20)
-According to [ghadimi2018approximation], we have
+According to [9], we have
 $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x)).$
@@
 Proof of Corollary˜5.3
-When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [jiang2024primal], we can find $\hat{v}$, $\hat{z}$ such that
+When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [16], we can find $\hat{v}$, $\hat{z}$ such that
 $\|v_{0}-v^{*}(x_{0})\|\leq\delta$
@@
 Appendix F Detailed Experimental Settings and Additional Experiments
-We adapt and modify the code from [jiang2024primal].
+We adapt and modify the code from [16].
 The experiments on the toy example and hyperparameter optimization for SVM are conducted on an AMD EPYC 9554 64-Core Processor.
@@
 To achieve this, slack variables $\xi$ are introduced to quantify classification violations for each sample.
-We consider the same problem formulation as in [jiang2024primal], which is
+We consider the same problem formulation as in [16], which is
 $\min_{c}\quad\mathcal{L}_{\mathcal{D}_{\text{val}}}(w^{*},b^{*})=\sum_{(z_{\text{val}},l_{\text{val}})\in\mathcal{D}_{\text{val}}}\exp\left(1-l_{\text{val}}\left(z_{\text{val}}^{\top}w^{*}+b^{*}\right)\right)+\frac{1}{2}\|c\|^{2}$
@@
 Then upper level variables are $c$, and the lower level variables are $w,b$. Thus, we have a coupled linear constraint for the lower level problem.
-We compared our algorithm SFLCB with GAM [xu2023efficient], LV-HBA [yao2024constrained], BLOCC [jiang2024primal], and BiC-GAFFA [yao2024overcoming] on the diabetes dataset.
-For GAM, we follow the same implementation approach and hyperparameters as [jiang2024primal], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [xu2023efficient].
+We compared our algorithm SFLCB with GAM [46], LV-HBA [50], BLOCC [16], and BiC-GAFFA [49] on the diabetes dataset.
+For GAM, we follow the same implementation approach and hyperparameters as [16], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [46].
 For BLOCC and SFLCB, we set the LL objective as $\frac{1}{2}\|w\|^{2}+\frac{\mu}{2}b^{2}$, where $\mu=0.01$ serves as a regularization term to make the LL problem strongly convex w.r.t $w,b$.
@@
 For our algorithm SFLCB, we set $\delta=0.01,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.001,\rho_{1}=\rho_{2}=0.01$.
-The hyperparameters of LV-HBA, BLOCC are the same as those used in [jiang2024primal] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [yao2024overcoming].
+The hyperparameters of LV-HBA, BLOCC are the same as those used in [16] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [49].
 The experiments are conducted across 10 different random train-validation-test splits, and the average results along with one standard deviation are reported in Figure˜2.
 F.2 Transportation network design
-We consider the same setting as in [jiang2024primal].
+We consider the same setting as in [16].
 In this setting, we act as the operator to design a new network that connects a set of stations $S$.
@@
 The fraction of passengers choosing the new network for each $(o,d)$ pair is denoted by $y_{od}$, and $y^{ij}_{od}$ represents the proportion using link $(i,j)$.
-We summarize the notation as follows. We keep the notation the same as in [jiang2024primal].
+We summarize the notation as follows. We keep the notation the same as in [16].
   • $x_{ij}\in\mathbb{R}_{+}$, the capacity of the new network for the link $(i,j)\in\mathcal{A}$.
@@
 where (47) are the flow-conservation constraints and (48) are the capacity constraints.
-We consider the same 3 nodes and 9 nodes settings as in [jiang2024primal] and use the same lower level feasibility criteria.
+We consider the same 3 nodes and 9 nodes settings as in [16] and use the same lower level feasibility criteria.
 The hyperparameter settings used in Figure˜3 are listed below.
@@
 For the 9 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.25$, $\rho_{1}=\rho_{2}=50$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-5$, and $T=300000$.
-For BLOCC, we used the same hyperparameters as those used in [jiang2024primal].
+For BLOCC, we used the same hyperparameters as those used in [16].